### v1

# Fight like a Woman:

Abstract
This article investigates the impact of judges’ gender on the outcome of domestic violence cases. Using data From São Paulo, Brazil, between 2010 and 2019, we compare conviction rates by judge’s gender and find that a domestic violence case assigned to a female judge is 31% (10 p.p) more likely to result in conviction than a case assigned to a male judge with similar career characteristics. To show that this decision gap rises due to different gender perspectives about domestic violence instead of rising due to female judges being tougher than male judges, we compare it against gender conviction rate gaps in similar types of crimes. We find that the gender conviction rate gap for domestic violence cases is significantly larger than the same gap for other misdemeanor cases (3 p.p. larger) and for other physical assault cases (8 p.p. larger). Lastly, we find evidence that at least two channels explain this gender conviction rate gap for domestic violence cases: gender-based differences in evidence interpretation and gender-based sentencing criteria.
Keywords: Domestic Violence, Judicial Decisions, Gender Bias
JEL Codes: D73, J16, K42
Introduction
Violence against women is frequently perpetrated inside the home by the ones closest to them. The United Nations Office on Drugs and Crime (2023) estimates that 55% of all homicides against women worldwide in 2022 were committed by intimate partners or other family members. As this violence is a criminal offense, a key component of the fight against domestic violence is the Criminal Court System and, therefore, how court judges approach this type of crime (Beecher-Monas, 2001). For instance, the HumanRightsWatch (1991) argues that, in the 1980s and early 1990s, Brazilian judges failed to analyze crimes of domestic violence against women in a non-discriminatory manner. [^1] In the years since, Brazilian authorities have made important legal and institutional changes to address domestic violence crimes. The most prominent of these is the Maria da Penha Law, enacted in 2006, which characterizes domestic violence as a specific type of criminal offense. Moreover, Kafka et al. (2019) finds that American judges in North Carolina use gender stereotypes to argue that some women want to “game the system” with “frivolous cases” when requesting protective orders against their abusive partners.
Since discriminatory gender stereotypes may vary depending on the decision maker’s individual characteristics, we investigate the impact of judges’ subjectivity on the outcome of domestic violence cases. In particular, we focus on the effect of a judge’s gender on their decisions, documenting a gender conviction-rate gap specific to domestic violence cases. We also analyze possible forces driving this difference.
To do this, we collect data from São Paulo, Brazil, between 2010 and 2019 and compare conviction rates by the deciding judge’s gender after controlling for the judge’s career characteristics. Case characteristics are independent of the presiding judge’s gender because cases are randomly allocated to judges within each court district in Brazil. Hence, the gender conviction-rate gap captures different decision criteria by judges of different genders.
In this analysis, we find that a domestic violence case assigned to a female judge is 31% (10 p.p) more likely to result in conviction than a case assigned to a male judge with similar career characteristics. Since this decision gap may be due to differences between male and female judges regardless of their perspectives on domestic violence, we compare this gap to gender decision gaps for two other types of crime: misdemeanors and physical assault. Because we focus on domestic violence crimes that do not result in death or long-term physical impairment, misdemeanors are comparable to domestic violence cases in their sentence severity (according to Brazilian Law). Physical assault cases are comparable to domestic violence cases because both involve some sort of physical aggression. We find that the gender conviction rate gap for domestic violence cases is more than twice the size of that of other misdemeanors. There are no differences between female and male judges in sentencing physical assault cases that do not involve domestic violence. Consequently, the gender conviction rate gap for domestic violence cases likely arises due to different perspectives about this type of crime, rather than because female judges are in general more punitive than their male counterparts.
We find evidence of two drivers of the gender conviction rate gap. The first is sharing an identity with the typical victim of domestic violence cases. To investigate if the salience of gender identity matters, we test if the gender difference in strictness is more pronounced in intimate partner violence cases. The incidence and tolerance of this type of offense is strongly associated with patriarchal gender norms involving partners (Heise and Kotsadam, 2015; González and Rodríguez-Planas, 2020). To do this, we take advantage of two key features of our dataset. First, judges are randomly allocated to cases, which makes the relationship between the victim and the assailant independent of judge characteristics. Second, we are able to identify the victim–assailant relationship in cases that result in conviction. We find a statistically significant increase (37.5%)in the likelihood of conviction in a case involving intimate partner violence if the judge is a woman. However, the probability of conviction in a case not involving intimate partner violence is the same for men and women judges.
Second, we test for gender differences in interpretation of the evidence provided in domestic violence cases. To test for the relevance of this driver, we compare the gender conviction rate between cases in which the defendant was caught by the police while committing the crime and cases without this objective piece of evidence. We find that male and female judges make similar decisions when analyzing cases where defendants are caught in the act, while judges differ when making decisions about cases with more subjective evidence. Hence, female judges interpret evidence differently from their male peers when analyzing domestic violence cases. This may mean that judges can make different decisions about what constitutes “true domestic violence” versus victims “gaming the system,” as characterized in Kafka et al. (2019), or that judges may disagree about the appropriate severity of punishment when there is room for discretion.
Finally, we document that the difference in conviction rates between male and female judges is more pronounced later in their careers, which is when career pressures are weaker. We define late-career judges as those in the last stage of the career of a Trial Court judge. Because judges in Brazil are not elected and have job security, [^2] By law, they acquire job stability after 2 years on the job. The goal is to insulate them from political pressure. promotion is the main form of career incentive. However, promotion from the Trial Court to the Appeals Court is rare and extremely unlikely in the short run, unlike career progression at earlier stages. We find that there is no gender conviction rate gap for junior judges, while senior female judges convict domestic violence cases more frequently than male judges in the same career stage. This is an important avenue for future research because it could mean that the difference in perspective of early-career judges is not entirely reflected in the gender conviction gap, potentially because of early-career judges’ desire to conform to the court’s established norms.
This article contributes to the literature about in-group bias in the judicial system. A large body of work finds evidence of judges making different decisions when they share a group identity with defendants or plaintiffs (Kruttschnitt and Savolainen, 2009; Shayo and Zussman, 2011a, 2017; Bielen and Grajzl, 2020; Cai et al., 2022; Chen et al., 2022). Similarly to our work, Kruttschnitt and Savolainen (2009), Bielen and Grajzl (2020), and Chen et al. (2022) analyze in-group gender bias in cases of uncommon crime types (rape and human trafficking), civil cases, and for female offenders, respectively. Our work differs from theirs in that we focus on a single common crime type (domestic violence) whose victims are mostly women.
The paper is organized as follows. Section 2 describes our dataset, while Section 3 explains our empirical strategy. In Section 4, we document the gender conviction rate gap, while 5 discusses its driving forces. Section 7 concludes.
Data Description and Summary Statistics
To analyze the judicial gender decision gap in domestic violence cases, we use data on criminal cases’ characteristics and the judges analyzing those cases. Section 2.1 summarizes our criminal case dataset, while Section 2.2 describes our dataset of judges deciding these cases. Lastly, Section 2.3 explains which judges are included in our sample and presents summary statistics.
Data on Criminal Cases
We collect data from all criminal cases brought to the Justice Court System in the State of São Paulo, Brazil, between January 4th, 2010, and December 3rd, 2019. This dataset includes the date each case was assigned to a courtroom, the court district where the crime took place, the judge assigned to the case, the full text of the sentence, a list of the main events of the case, and the primary and secondary crime types associated to the case.
The data also contain two pieces of information that are crucial for the construction of our sample, which we discuss in detail in Section 2.3. The first is a dummy indicating if the case was randomly assigned to a judge within a court district and point in time [^3] That is, not connected to other crime randomly assigned previously.. The second is the name of the courthouse, where different courthouses correspond to certain legal specializations. We use this information to make sure that domestic violence cases are indeed randomly assigned within a district and not systematically assigned to (or reassigned to) judges who do not preside over cases of other types of crimes.
We use the full-sentence text to construct our main outcomes. To do this, we adapt the classification algorithm developed by Possebom (2023) so that it classifies domestic violence case outcomes in a more precise manner. [^4] See Appendix A for an overview of our data-construction process. We classify sentences into one of six possible judicial decision categories.
The first decision type is conviction. This variable equals one when the judge convicts at least one of the defendants listed in the case of at least one of the charges on the case. As a consequence of conviction, the defendant receives some type of punishment (e.g., incarceration, fine, or community service) and a criminal record.
The second decision category is a non-prosecution agreement (NPA). This variable equals one when the defendant and the district attorney’s office strike a deal to exchange a criminal charge for a lighter punishment, such as a fine. Such agreements must be authorized by a judge and do not generate a criminal record.
The third decision type is dismissal. A judge may dismiss a case if they consider it a less serious offense and the defendant fulfills a series of requirements. For example, the defendant may be forced to show up in court periodically for two years and be forbidden from going to bars. Similarly to NPAs, dismissed cases do not generate criminal records.
The fourth decision category is acquittal. This variable equals one if the defendant is ruled “not guilty” on the charges listed.
Lastly, we account for cases that do not come to a conclusion because either the defendant died before sentencing or because the criminal charges expired. In the latter situation, defendants can no longer be convicted according to the statute of limitations because too much time has elapsed since the day the crime was committed.
We use the list of the case’s main events to measure two key case characteristics. First, we measure whether the defendant was caught in the act of committing the crime (in flagrante delicto), which is considered strong evidence against the defendant. Second, we observe whether the sheriff temporarily detained the defendant in jail before the case went to the district attorney’s office and was assigned a judge. Sheriffs, who must have a Law degree in Brazil, may temporarily detain a defendant if they believe that the defendant poses a significant threat to society.
The district attorney’s office defines each case’s primary and secondary crime prior to judge assignment. We construct three samples based on those crime types.
First, our ‘domestic violence’ sample contains all cases with (i) domestic violence as the main subject or (ii) physical assault as the primary type and domestic violence as the secondary type. The classification “domestic violence” is based on the Maria da Penha Law (Law n. 11,340/2006) and increases the punishment severity of physical assault crimes associated with domestic violence.
We define a ‘physical assault’ sample of cases whose primary type is physical assault. These cases are associated with other forms of assault such as street fights, bar fights, or football hooliganism. We call ‘other physical assault’ the physical assault cases that do not involve domestic violence.
Third, we define a ‘misdemeanors’ sample comprised of those crimes that carry a maximum incarceration sentence of less than four years. This threshold is relevant because, according to Brazilian Law, being convicted of these crime types leads to a criminal record, but not to incarceration. Sentences must be converted into fines or community service. As discussed by Possebom (2023) and Acerenza, Possebom, and Sant’Anna (2024), most misdemeanor crimes in Brazil are theft offenses. We call ‘other misdemeanors’ the misdemeanor offenses unrelated to domestic violence.
Note that the sample of domestic violence cases, other physical assault cases, and other misdemeanor cases are mutually exclusive. We present summary statistics for each of these sub-samples in Section 2.3.
Judges’ Career and Characteristics
We collect information on judges’ careers from two public sources of information: the annual seniority list published in the São Paulo Justice Diary (2007 to 2019) and the judges’ monthly productivity spreadsheet from the São Paulo Court of Justice website (2011 to 2019).
The seniority list ranks all active judges in the court system as of December 31st of each year by career stage and seniority. The dataset includes their starting date as magistrates and the date they were promoted to their current career stage. We combine these pooled seniority lists into a dataset on judges’ careers following Laneuville (2024). [^5] This dataset allows us to track all variations of judges’ names over time. Names may vary over time due to marriage, divorce, or data entry errors. The judge-level dataset includes the date they entered the justice system, the date of every promotion they have received since December 31st, 2007, and, if they entered the justice system before 2007, the date they reached the career stage they were in at the end of 2007. We link this dataset to the criminal cases dataset by matching the judge’s name.
The productivity data from the São Paulo Court of Justice website contains records on the total amount of work done by each judge in any given month and court district from January 2011 to December 2019.666For example, it measures the number of hearings or disposed cases per month for each judge in each court district. We use this information to keep track of which judges are active in each court district-month pair. [^7] Months are the natural unit of time for keeping track of activity and work volume because salaries in Brazil are paid monthly. Furthermore, most hiring, reallocation, and termination takes effect at the beginning of each month.
Sample Definition and Summary Statistics
Our final dataset is at the criminal-case level. It is crucial for our identification strategy that, for each case, the gender of the judge is randomly assigned and correctly identified.
First, we restrict the dataset to cases that were randomly allocated to judges. By Brazilian Law, a case can be allocated to judges by either connected distribution and free distribution. In the first procedure, the crime under consideration is connected to an existing criminal process. Here, the current case is then attached to the previous process and analyzed by the judge associated with the older case. [^8] For two simple examples, a case where a murderer then stole a car to flee the police or a case where a drug trafficker was then caught laundering money. In the second procedure, the crime under consideration is not connected to an existing criminal process. The new case is then allocated to a judge working in the court district where the offense occurred. If there is more than one judge working in that court district, Brazilian Law dictates that the case be randomly allocated to any judge in the court district. To ensure that the cases in our sample are randomly allocated to judges (and therefore uncorrelated with judge gender), we focus on cases that are allocated according to the free distribution system.
Second, since we focus on domestic violence cases, we must guarantee that this type of case is randomly allocated to a judge. This assumption would be violated if a specific judge or courtroom within the court specialized in domestic violence cases. If there is such a specialized courtroom, then domestic violence cases bypass the free distribution system and are allocated to the specialized judge. For this reason, we restrict the sample to district-quarter pairs without a courtroom specializing in domestic violence.
Third, since we compare decisions made by female and male judges, we require the judge’s gender to be randomly allocated to cases. In other words, the ex-ante probability of a female judge being assigned to each case must be different from zero or one, which would not be the case in courts without judges of both genders. To enforce this restriction, we limit the sample to cases from court district-month pairs with at least one male and one female active judge in the productivity dataset.
Last, we need to ensure that the assigned judge’s gender is determined correctly. To do this, we focus on cases were allocated to judges between January 2011 and December 2019 and use our productivity dataset to remove all observations in which the assigned judge was not active at the court district and month of the assignment. [^9] Cases are randomly allocated to courtrooms and instead of directly to individual judges. If a court district is understaffed at the time of case assignment and there is no presiding judge in the assigned courtroom assigned, the system may record the case as being attributed to the next person who takes over the courtroom, which may take place at a later point in time.
Using this final sample, we compute summary statistics for each possible case outcome and crime type. Table 1, displays case outcomes in the columns and crime types in the rows. Domestic violence cases are more frequently acquitted than other misdemeanors or physical assault cases. As conviction rates are similar across types of crimes, this higher acquittal is due to fewer NPAs and dismissed cases.
Table 1: Case Outcomes by Crime Type in %
Crime Type Conviction NPA Acquittal Dismissal Expired Dead # Cases
(1) (2) (3) (4) (5) (6)
DV 32.5 1.1 51.4 4.3 8.8 2.0 4,358
Other Mis 36.9 5.4 31.2 15.5 9.5 2.2 39,815
Other PA 27.2 7.0 34.1 20.6 10.4 1.9 1,648
  • Notes: The rows of this table indicate the different samples used in this paper: domestic violence cases (DV), misdemeanor cases unrelated to domestic violence (Mis), and physical assault cases that are not domestic violence (PA). The first six columns indicate the frequency of each type of case outcome. The conviction dummy equals one when at least one defendant is convicted in at least one charge. The NPA dummy equals one when the defendant and the district attorney’s office sign an NPA. The dismissal dummy equals one if the case is dismissed. The acquittal dummy equals one if the defendant is found not guilty. The “expired” dummy equals one if at least one of the charges in the case has expired. The “Dead” dummy equals one if the defendant died the sentence’s date. The rows do not sum to 100% because cases with multiple charges can have different outcomes for each charge.
Empirical Strategy and Validity Tests
In this section, we explain the empirical strategy used to measure the gender conviction rate gap and indirectly test the strategy’s validity.
We measure the gender conviction rate gap for domestic violence cases by comparing the conviction probability for similar cases allocated to female and male judges. We do this by running a linear regression model on the sample of domestic violence crimes:
$\text{Convicted}_{it}=\beta\cdot\text{Female Judge}_{it}+\eta\cdot X_{it}+\text{Court District}_{i}\times\text{Quarter}_{t}+\varepsilon_{it}$ (1)
where $\text{Convicted}_{it}$ indicates whether the cases resulted in a conviction, $\text{Female Judge}_{it}$ indicates whether the case was assigned to a female judge, $X_{it}$ are covariates including case characteristics (an indicator for whether the defendant was already in jail and an indicator for whether the police caught them committing the crime) and pre-determined judge characteristics (career stage, total experience, and experience in their current career stage as of the day the case was assigned), and $\text{Court District}_{i}\times\text{Quarter}_{t}$ is a full set of court district-quarter pair fixed effects.
Our parameter of interest, $\beta$, measures whether a female judge is more likely to convict than a male judge under the assumption that the judge’s gender is independent of the case’s potential outcome. This exogeneity assumption is valid according to Brazilian Law mandating cases are randomly allocated to judges within court districts (Section 2.3). We indirectly test the validity of this assumption in Table 2.
Even when our exogeneity assumption is valid, the $\beta$ parameter in Equation (1) may capture differences between male and female judges that are unrelated to their views on domestic violence. To measure whether this conviction rate gap is due to different gender perspectives about this type of crime, we compare this domestic violence conviction gap against the gap for similar crime types. Our sample therefore consists of domestic violence cases and one type of comparable criminal cases. We estimate the following linear regression model:
$\text{Convicted}_{it}$ $=\beta_{1}\cdot\text{Female Judge}_{it}+\beta_{2}\cdot\text{Domestic Violence}_{it}$
$\hskip 20.0pt+\beta_{3}\cdot\text{Female Judge}_{it}\times\text{Domestic Violence}_{it}$ (2)
$\hskip 20.0pt+\eta\cdot X_{it}+\text{Court District}_{i}\times\text{Quarter}_{t}+\varepsilon_{it},$
where $\text{Domestic Violence}_{it}$ indicates whether the case is associated with domestic violence. We use two categories of comparable crime types: (i) offenses with similar sentences (misdemeanors) and (ii) offenses of a similar nature (physical assault).
In Equation (2), our parameter of interest is $\beta_{3}$. This specification accounts for the extra strictness and social unacceptability of domestic violence offenses beyond that of other violent crimes of comparable severity. Consequently, this coefficient captures the part of the gender conviction rate gap that is likely due to different gender perspectives about domestic violence cases.
This interpretation for the $\beta_{3}$ coefficient is valid if the judge’s gender is independent of the case’s potential outcome after conditioning on the type of crime. This exogeneity assumption is again valid according to Brazilian Law since cases are randomly allocated to judges within court districts and crime types are determined by the district attorney prior to judge assignment (Section 2.3). We again test this assumption’s validity in Table 2.
To test the validity of the identifying assumptions behind Equations (1) and (2), we measure the linear relationship between case characteristics and a dummy variable for an assigned female judge. Table 2 shows these correlations for each sample of crimes. We find that, after including court district-quarter fixed effects, judge gender is uncorrelated with case characteristics such as whether the defendant was in jail before the case was assigned, whether the defendant was caught in the act by the police and whether the crime is a domestic violence offense. This result suggests that our identifying assumptions are valid.
Table 2: Testing the Identification Assumption: Gender of the Judge and Case Characteristics
Female Judge Female Judge Female Judge
(1) (2) (3)
Domestic Violence $0.01$ $-0.01$
$(0.02)$ $(0.01)$
Caught red-handed? $0.01$ $0.00$ $0.02$
$(0.03)$ $(0.01)$ $(0.03)$
Defendant in Jail $-0.02$ $-0.01$ $-0.02$
$(0.03)$ $(0.01)$ $(0.03)$
Sample Physical Assault Misdemeanors Domestic Violence
Quarter x Court District FE YES YES YES
Num. obs. $6,006$ $44,173$ $4,358$
${}^{***}p<0.01$; ${}^{**}p<0.05$; ${}^{*}p<0.1$
  • Notes: Each cell represents the linear relationship between the gender of the judge randomly assigned to the case and a set of characteristics of the case. Column 1 shows this relationship for the sample of all physical assault crimes, Column 2 shows this relationship for the sample of misdemeanors (crimes such that the maximum sentence does not exceed 4 years). Column 3 shows this relationship for the sample of domestic violence crimes. All regressions include quarter-by-court district fixed effects.
Measuring the Gender Conviction Rate Gap
We measure the gender conviction rate gap using two empirical strategies. First, we estimate Equation (1) and find that female judges are more likely to convict in domestic violence cases than equally senior men. Second, we estimate Equation (2) and find that the gender conviction rate gap is small or nonexistent for misdemeanors and non-domestic violence physical assaults, whereas the gap becomes large and significant for domestic violence cases.
Table 3 shows the effect of judge gender on the probability of conviction in domestic violence cases according to Equation (1). In the first column, we find that randomly selecting a woman or a man to preside does significantly impact conviction rates.
However, as shown in Laneuville (2024), the average male judge and the average female judge have different career trajectories. We may therefore be concerned that judge tenure and experience may have an impact on conviction rates. For this reason, Column (2) controls for judge career stage, number of years of experience, and number of years in their current position. Here, we find a large effect of gender on the probability of conviction: a woman judge is 31.2% (10 p.p.) more likely to convict a domestic violence offender than a male judge with similar experience. Column (3) shows that this result is robust to the introduction of other case controls.
Table 3: Effect of Judge Gender on the Probability of Conviction in Domestic Violence Cases
Convicted Convicted Convicted
(1) (2) (3)
Fem $0.04$ $0.10^{***}$ $0.10^{***}$
$(0.03)$ $(0.03)$ $(0.03)$
Experience $0.01^{***}$ $0.01^{***}$
$(0.00)$ $(0.00)$
Years in Career Stage $-0.01^{*}$ $-0.01^{*}$
$(0.01)$ $(0.01)$
Caught Red Handed? $0.06$
$(0.04)$
Defendant in jail $-0.03$
$(0.04)$
Male Judge’s Conviction Rate $0.32$ $0.32$ $0.32$
Quarter x Court District FE YES YES YES
Career Stage Controls NO YES YES
Num. obs. $4,358$ $4,358$ $4,358$
${}^{***}p<0.01$; ${}^{**}p<0.05$; ${}^{*}p<0.1$
  • Notes: This table presents the estimated results of Equation (1). The first column shows the effect of judge gender on the probability of conviction. The second column includes controls for judge characteristics (career stage, number of years of experience as a judge, and number of years of experience in the current career stage). The third column controls for both judge and case characteristics. All regressions include quarter-by-court district fixed effects.
Since the results in Table 3 may capture differences between male and female judges beyond their gendered perspectives on domestic violence, we also estimate Equation (2) to compare the gender conviction rate gap for domestic violence cases against the gap for similar cases. Table 4 shows the estimated results associated with Equation (2). The first three columns use the sample of misdemeanors, therefore comparing crimes with similar sentence lengths to those in the domestic violence sub-sample. The last three columns use the sample of physical assault cases, which are similar in nature to domestic violence cases as they entail physical harm to another person.
Table 4: Effect of Judge Gender on the Probability of Conviction Comparing Domestic Violence Cases against Similar Crimes
Sample: Misdemeanors Sample: Physical Assault
Convicted Convicted Convicted Convicted Convicted Convicted
(1) (2) (3) (4) (5) (6)
Female $0.02^{**}$ $0.02^{***}$ $0.02^{***}$ $-0.03$ $-0.00$ $0.00$
$(0.01)$ $(0.01)$ $(0.01)$ $(0.04)$ $(0.04)$ $(0.04)$
DV $-0.05^{***}$ $-0.05^{***}$ $-0.04^{***}$ $0.06^{***}$ $0.06^{**}$ $0.05^{**}$
$(0.01)$ $(0.01)$ $(0.01)$ $(0.02)$ $(0.02)$ $(0.02)$
DV x Female $0.03^{*}$ $0.04^{*}$ $0.03^{*}$ $0.07^{*}$ $0.09^{**}$ $0.08^{**}$
$(0.02)$ $(0.02)$ $(0.02)$ $(0.04)$ $(0.04)$ $(0.04)$
Men’s CR (Non-DV) $0.38$ $0.38$ $0.38$ $0.30$ $0.30$ $0.30$
Men’s CR (All) $0.37$ $0.37$ $0.37$ $0.31$ $0.31$ $0.31$
Career Controls NO YES YES NO YES YES
Case Controls NO NO YES NO NO YES
Quarter-District FE YES YES YES YES YES YES
Num. obs. $44,173$ $44,173$ $44,173$ $6,006$ $6,006$ $6,006$
${}^{***}p<0.01$; ${}^{**}p<0.05$; ${}^{*}p<0.1$
  • Notes: This table presents the estimated results of Equation (2). The first three columns compare differences in male and female judge conviction rates in domestic violence cases to the difference in conviction rates for misdemeanors. The last three columns compare differences in male and female judge conviction rates in domestic violence cases to the difference in conviction rates for other physical assault crimes. CR stands for Conviction Rate and DV stands for Domestic Violence.
Table 4 shows that the estimated coefficient on the female dummy ($\beta_{1}$) is small for the misdemeanors sample and close to zero for the physical assault sample. This suggests that the gender conviction rate gap is small or nonexistent for comparable crimes that are not classified as involving domestic violence.
More importantly, Table 4 shows that the coefficient on the interaction of judge gender and domestic violence ($\beta_{3}$) is statistically significant and positive in all regressions, regardless of included control variables. The effect is particularly strong for the physical assault sample (Columns (4)-(6)). This result suggests that the difference in conviction rates verified in Table 3 is neither fully explained by the severity or the nature of domestic violence offenses nor by female judges being tougher than male judges across the board.
The interaction coefficient captures the part of the gender conviction rate gap that is specific to domestic violence cases and is likely due to different approaches to domestic violence cases between male and female judges. In the next section, we explore this difference more deeply.
Understanding the Gender Conviction Rate Gap
In this section, we show two important findings about the conviction rate gender gap. The first is suggestive evidence that the gap is explained by judges’ identification with their own gender. We find that the difference in conviction rates between male and female judges is more intense in domestic violence cases involving intimate partner violence. Violence between relationship partners is highly impacted by patriarchal social norms that make gender identity more salient.
The second finding, discussed in Section 5.2, is that men and women respond differently to evidence presented in the same case. We present evidence that female judges are more likely than male judges to convict in cases where evidence is more tenuous and subject to interpretation, but not when the evidence is objectively incriminating.
Female Identity as a Driver for Conviction Rates
Next, we show that the gender of the judge matters more in cases where patriarchal social norms are particularly relevant: intimate partner violence. We interpret this finding as indicative that a major reason for the gender conviction gap is that domestic violence cases make gender identity more salient to judges.
First, sexist social norms and gender roles significantly affect the occurrence and punishment of intimate partner violence in multiple contexts. For instance, González and Rodríguez-Planas (2020) show that, among immigrant families, the incidence of intimate partner violence is positively affected by gender norms in the country of origin. Heise and Kotsadam (2015) show that the incidence of intimate partner violence in different countries is strongly associated with patriarchal social norms, such as male authority over women, justifications for wife beating, and laws and practices that disadvantage women in property ownership and other economic resources. Closer to our context, Perova and Reynolds (2017) show that police stations that specialize in domestic violence with specially trained officers significantly reduce the number of deaths of young women by intimate partner violence in metropolitan municipalities in Brazil. In terms of judicial decision-making on domestic violence, the negative stereotypes referred to in Kafka et al. (2019) and HumanRightsWatch (1991) concern female intimate partners.
We then take advantage of an important feature of our data: we are able to identify the relationship between the victim and the assailant for all cases that culminate in conviction. [^10] In fact, we observe this relationship for cases in which the judge must make an active decision, e.g., excluding cases where the judge declares the defendant dead, a non-prosecution agreement is reached, or the case expires. To do this, we construct an algorithm that defines a case as intimate partner violence if, when writing the court’s decision, the judge uses the Portuguese word for wife, girlfriend, ex-wife, ex-girlfriend, or similar words. [^11] In the domestic violence sample, 57.8% of cases are flagged as violence against partners or ex-partners. The specific words we look for are “namorada,” “ex-namorada,” “esposa,” “ex-esposa,” “amásia,” “ex-amásia,” “companheira,” “ex-companheira,” “mulher,”, “ex-mulher,” “coabitação,” “união estável,” “casamento,” “era casado,” and “era casada.”
To show that the gender of the judge matters more for the outcome of intimate partner violence cases than for that of other domestic violence offenses, we make the following argument. If the probability of a conviction for non-intimate partner violence is non-zero, then two things cannot be true at the same time: 1) the probability of conviction for a domestic violence offense involving intimate partner violence is higher for women than it is for men; 2) the probability of conviction for a domestic violence offense not involving an intimate partner is not higher for women than it is for men. Although we cannot observe the relationship between victim and assailant in all cases, we can reasonably assume that this relationship is independent of the assigned judge’s gender due to the fact that cases are randomly allocated to judges in our sample.
To test for this possibility, we interact the indicator for assault between intimate partners with the conviction indicator. This interaction term becomes our new outcome variable in Equation (1).
Table 5 shows the results of these regressions. In Column (1), the outcome is a dummy variable indicating whether the case included intimate partner violence and resulted in a conviction. In Column (2), the outcome is a dummy variable indicating whether the case did not include intimate partner violence and also resulted in a conviction. In Column (3), the outcome is a dummy variable indicating whether the case resulted in a conviction, replicating the results from Table 3 for convenience.
We find that the probability of an intimate partner violence case resulting in conviction is 37.5% (9 p.p.) higher when the judge is female (Column (1)). However, the probability of a case not involving intimate partner violence resulting in conviction is the same regardless of the gender of the judge (Column (2)). These results suggest that the gender conviction rate gap is higher in cases where a female judge may identify more strongly with the victim compared to a male judge.
Table 5: Effect of Judge Gender on the Probability of Conviction and the Salience of Intimate Partner Violence
Convicted and Convicted and Convicted
Partner Not Partner
(1) (2) (3)
Female $0.09^{***}$ $0.01$ $0.10^{***}$
$(0.03)$ $(0.01)$ $(0.03)$
Men’s Outcome Average $0.24$ $0.07$ $0.32$
Quarter-District FE YES YES YES
Case Controls YES YES YES
Judge Career Controls YES YES YES
Num. obs. $4,358$ $4,358$ $4,358$
${}^{***}p<0.01$; ${}^{**}p<0.05$; ${}^{*}p<0.1$
  • The regressions in this table refer to outcomes for the sample of domestic violence crimes. The outcome in Column (1) is whether the case involved intimate partner violence and led to a conviction. The outcome in Column (2) is whether the case did not involve intimate partner violence and led to a conviction. The outcome in Column (3) is whether the case led to a conviction. All regressions include quarter-district fixed effects and a full set of controls for case characteristics and judge career stages.
These findings are consistent with existing evidence that, in many contexts, judges make different decisions when they share a group identity with defendants or plaintiffs (Kruttschnitt and Savolainen, 2009; Shayo and Zussman, 2011a, 2017; Bielen and Grajzl, 2020; Cai et al., 2022; Chen et al., 2022). Many of these papers show that the differences intensify as group identity becomes more salient. This intensification may result from of a heightened opposition between two or more social groups (“us” versus “them”). For example, Shayo and Zussman (2011b) show that Jewish Israeli judges in small claim courts favor Jewish litigants more intensely when they are more exposed to conflict-related terrorism. Closer to our case, Cai et al. (2022) show that judges analyzing divorce cases in China favor plaintiffs of their own gender more in areas where social norms are more conservative. In our case, we show that the gender gap is much more intense in cases where the form of domestic violence is more stereotypical [^12] Note that most domestic violence against women is committed by partners and ex-partners (United Nations Office on Drugs and Crime, 2023). and has the potential to cause the most division because of social norms. Judges’ views on other forms of domestic violence may be less dependent on their views on patriarchy. For example, beating one’s own parents or children is more likely to be viewed as unjustifiable by men and women alike.
Male and Female Judges Analyze Evidence Differently
Domestic violence cases require complex evidentiary rulings. As discussed by Beecher-Monas (2001), these cases may involve both victim testimony about the incident(s) and expert witness testimony about battered woman syndrome. Since such testimony is inherently subjective, female and male judges may interpret testimony evidence differently.
Our dataset allows us to investigate whether the gender conviction rate gap is stronger in cases where the evidence set is less clearly incriminating. For each case, we observe whether the defendant was arrested because they were caught in the act of committing the crime (i.e., in flagrante delicto). These cases have more objective evidence, often including testimony from police officers who witnessed the offense.
If the gender conviction rate gap is stronger in cases where the defendant was not caught red-handed, then this may be because female and male judges interpret more subjective pieces of evidence differently. This would provide possible explanatory force for the observed leniency gap in domestic violence cases. If the gender conviction rate gap is similar across domestic violence cases regardless of the type of evidence presented, then there must be other forces driving this lenience gap.
To measure the gender conviction rate gap for cases with different evidence types, we re-estimate Equations (1) and (2) using two sub-samples. The first restricts our sample to cases where the defendant was caught red-handed, while the second is restricted to cases where the defendant was not caught in flagrante delicto.
Table 6 reports regression results for the two different samples. The odd columns include cases where the defendant was caught red-handed, while the even columns include cases where this objective evidence was not presented. Columns (1) and (2) report the results associated with Equation (1), while Columns (3)-(6) report the results associated with the interaction model described in Equation (2).
Table 6: Effect of Judge Gender on the Probability of Conviction conditional on the Type of Evidence
Sample: Sample: Sample:
Domestic Violence Misdemeanors Physical Assault
In flagrante delicto Yes No Yes No Yes No
(1) (2) (3) (4) (5) (6)
Fem $0.04$ $0.13^{***}$ $0.02$ $0.03^{**}$ $0.05$ $0.01$
$(0.06)$ $(0.04)$ $(0.01)$ $(0.01)$ $(0.13)$ $(0.04)$
DV $-0.14^{***}$ $-0.01$ $0.03$ $0.07^{**}$
$(0.02)$ $(0.01)$ $(0.08)$ $(0.03)$
DV x Fem $0.05$ $0.03$ $-0.02$ $0.12^{**}$
$(0.03)$ $(0.02)$ $(0.13)$ $(0.05)$
Men’s CR (Non-DV) - - $0.49$ $0.33$ $0.40$ $0.28$
Men’s CR (All) $0.35$ $0.30$ $0.47$ $0.32$ $0.36$ $0.30$
Quarter-District FE YES YES YES YES YES YES
Judge Career Controls YES YES YES YES YES YES
Num. obs. $1,337$ $3,021$ $14,391$ $29,782$ $1,552$ $4,454$
${}^{***}p<0.01$; ${}^{**}p<0.05$; ${}^{*}p<0.1$
  • Notes: This table compares the determinants of the probability of conviction. Cases are split into a sub-sample in which the defendant was caught red-handed and a sample in which the defendant was not, requiring the justice system to rely on less objective evidence. The first two columns are constructed using the sample of domestic violence cases. Columns 3 and 4 use the sample of misdemeanors. Columns 5 and 6 use the sample of physical assault crimes. CR stands for Conviction Rate and DV stands for Domestic Violence.
We find that a gender conviction rate gap exists only in those cases where the defendant was not caught red-handed. We only find significant results for the female coefficient ($\beta$) in Column (2) and the interaction coefficient ($\beta_{3}$) in Column (6). These results suggest that female and male judges interpret less objective evidence differently, while they interpret objective evidence (cases flagged as in flagrante delicto) in similar ways. These different interpretations of more subjective cases offer one explanation for the gender conviction rate gap in domestic violence cases.
Discussion: Career Incentives and a Potential Underestimation of the Gap
Next, we analyze the association between the strength of career incentives and the size of the conviction gender gap. We show that male and female judges differ more intensely at the end of their careers. This corresponds with the period where they face the lowest probability of upcoming performance evaluation for promotion to the Appeals Court. [^13] See Appendix B for more information on the proportion of cases analyzed by judges who apply for promotion in the near future Career pressures may lead early-career male and female judges to act more similarly, possibly in response to feedback from Appeals Court magistrates, who are overwhelmingly male.
Since Brazilian judges cannot be fired and must receive the same salary, promotion is the main career incentive faced by magistrates in our sample. Judges are unelected officials who may occupy four positions during their careers in the Trial Court (substitute, initial, intermediary, and final). At the final stage of their careers, they may be promoted to the Appeals Court. To be promoted to the next career stage, Trial Court Judges must apply for promotion opportunities and be approved by a vote of the Appeals Court Judges. Since very few judges are promoted to the Appeals Court (Laneuville, 2024), career incentives are stronger for trial judges who are not at the final stage of the Trial Court. [^14] Appendix B provides descriptive statistics showing that final-stage judges face weaker career incentives.
Since adopting different sentencing criteria from peer judges may hamper chances for promotion, we investigate whether judges who face weak career incentives are more likely to differ from their colleagues, and whether these differences correspond with judge gender. If we observe a larger gender conviction gap for final-stage judges, then weaker career incentives may explain the gender leniency gap in domestic violence cases documented in Section 4. If judges in all career stages have similar gender conviction rates, then there must be other forces driving this lenience gap.
To measure the gender conviction gap for judges in different career stages, we re-estimate Equations (1) and (2) using two sub-samples. The first restricts our sample to cases assigned to final-stage judges, while the second restricts the sample to cases assigned to judges in earlier stages of the Trial Court.
Table 7 presents the estimated coefficients of Equations (1) and (2) for these two sub-samples. Results for the sample of final-stage judges are provided in the odd columns. Results for the sample of early-stage judges are provided in the even columns. The header indicates whether the sample contains only domestic violence cases, misdemeanor cases, or physical assault cases.
Table 7: Effect of Judge Gender on the Probability of Conviction by Judge Career Stage
Sample: Sample: Sample:
Domestic Violence Misdemeanors Physical Assault
Final Early Final Early Final Early
(1) (2) (3) (4) (5) (6)
Fem $0.23^{***}$ $-0.03$ $0.06^{***}$ $-0.00$ $0.09^{**}$ $-0.13^{*}$
$(0.04)$ $(0.07)$ $(0.01)$ $(0.02)$ $(0.05)$ $(0.07)$
DV $-0.08^{***}$ $0.02$ $0.04$ $0.06$
$(0.01)$ $(0.02)$ $(0.03)$ $(0.05)$
DV x Fem $0.09^{***}$ $-0.02$ $0.10^{*}$ $0.11$
$(0.03)$ $(0.03)$ $(0.05)$ $(0.07)$
Men’s CR (Non-DV) - - 0.36 0.41 0.24 0.37
Men’s CR (All) $0.26$ $0.43$ $0.35$ $0.50$ $0.25$ $0.50$
Quarter-District FE YES YES YES YES YES YES
Case Controls YES YES YES YES YES YES
Num. obs. $2,609$ $1,749$ $25,639$ $18,534$ $3,499$ $2,507$
${}^{***}p<0.01$; ${}^{**}p<0.05$; ${}^{*}p<0.1$
  • Notes: This table shows results of regressing the conviction indicator on the indicator variables for female judges, domestic violence cases, and the interaction female x domestic violence for six different samples. Column (1) shows the results for the domestic violence cases assigned to final-stage judges. Column (2) shows the results for domestic violence cases assigned to early-stage judges. Column (3) shows the results for misdemeanor cases assigned to final-stage judges. Column (4) shows the results for misdemeanor cases assigned to early-stage judges. Column (5) shows the results for physical assault cases assigned to final-stage judges. Column (6) shows the results for cases assigned to early-stage judges. All regressions include quarter x court district fixed effects. CR stands for Conviction Rate and DV stands for Domestic Violence.
We find that the gender conviction rate gap is much stronger for final-stage judges than for early-stage judges. In the sample of domestic violence cases, final-stage female judges are almost twice as likely to convict as their male peers (23 p.p. larger), while early-stage judges have similar conviction rates across genders. We reach qualitatively similar conclusions when analyzing the interaction coefficients in the misdemeanors and other physical assault samples. Hence, the bulk of the gender gap discussed in this paper comes from judges at the end of their careers who face lower pressure to perform. If career pressures affect judges’ verdicts, then early-career judges of different genders may have different views on domestic violence than their conviction rates might suggest.
Conclusion
In this paper, we provide evidence that cases assigned to female judges are significantly more likely to result in conviction than those assigned to a male judges with a similar career path. This pattern is, to a large extent, due to the reasons that are specific to domestic violence. The gender difference in conviction rates is stronger for cases involving domestic violence than for crimes that are similar in nature (physical assault) and sentence length (misdemeanors).
Furthermore, we find suggestive evidence for two major drivers of the conviction gender gap in domestic violence. The first is that female judges are stricter than men in cases of intimate partner violence but equally strict for cases of other domestic violence offenses. We claim this is evidence of identification with one’s own gender, as perception of intimate partner violence is highly affected by patriarchal social norms, as opposed to other forms of violence where this might not be the case.
The second driver comes from gender differences in the way judges analyze evidence. There is no gender gap when the defendant is caught red-handed by the police, making the evidence presented clear and objective, but the gap is very strong when there is more room for subjectivity in the analysis.
Finally, we identify an association between a larger gender gap and weaker career pressures for more senior judges in the final stage of the trial court. We believe this to be an important avenue for future research.
Supporting Information
(Online Appendix)
Appendix A Additional Details about the Criminal Cases Data
In this appendix, we explain how we construct our dataset about criminal cases in São Paulo. We build upon the dataset collected by Possebom (2023), modifying it to allow for a deeper analysis of domestic violence cases. For additional legal definitions, check Appendix I.2 by Possebom (2023).
The final dataset was created from two initial datasets.
  1. 1.
    CPOPG (“Consulta de Processos de Primeiro Grau”): It contains information about all criminal cases in the Justice Court System in the State of São Paulo (TJ-SP) between 2010 and 2019.
  2. 2.
    CJPG (“Consulta de Julgados de Primeiro Grau”): It contains information about the last decision made by a trial judge in all criminal cases in TJ-SP between 2010 and 2019.
Starting from the CPOPG dataset, we implement the following steps.
  1. 1.
    We only keep cases that are currently in the Appeals Court, closed, or whose status is empty. Those cases are already associated with a trial judge’s sentence.
  2. 2.
    We only keep cases that aim to analyze whether a defendant is guilty or not.
  3. 3.
    We only keep cases that were randomly assigned to trial judges.
  4. 4.
    We only keep cases whose starting date is after January 1st, 2010.
We then merged it with the CJPG dataset using cases’ id codes.
After this step, we randomly select 65 cases per year (2010-2019) for manual classification. From this sample, we select 35 cases each year to form our training sample, ensuring that at least 10 of them are physical assault cases. Moreover, we select 10 cases each year to form our overall validation sample, and 10 domestic violence cases and 10 extra physical assault cases to form our crime-type-specific validation samples.
We manually classify these cases into six categories: “defendant died during the trial”, “case was expired”, “defendant is guilty”, “defendant accepted a non-prosecution agreement” (“transação penal” in Portuguese), “case was dismissed” (“processo suspenso” in Portuguese) and “defendant was acquitted”. Since some sentences are missing or incomplete, we are able to manually classify only 611 sentences.
Now, we use those 611 manually classified cases to train a classification algorithm. To do so, we divide them into a training sample (315 cases), an overall validation sample (109 sentences), a domestic violence-specific validation sample (99 sentences), and a physical assault-specific validation sample (88 sentences).
First, we design an algorithm to identify which defendants died during the trial. To do so, we check whether the sentence contains any reference to the first paragraph of Article 107 from the Brazilian Criminal Code. This deterministic algorithm classifies cases into the category “defendant died during the trial” almost perfectly. It misclassified only one sentence that had two defendants and only one of them died before the judge made a decision.
Second, we design an algorithm to identify which cases expired. To do so, we check whether the sentence contains any reference to the fourth paragraph of Article 107 from the Brazilian Criminal Code. This deterministic algorithm classifies cases into the category “expired case” almost perfectly. It misclassified only four sentences that included multiple crime types in one single case.
Third, we design an algorithm to identify which cases were dismissed. To do so, we check whether the sentence contains any reference to Article 89 in Law n. 9099/95. This deterministic algorithm correctly classifies 98% of the cases into the category “case was dismissed”.
Fourth, we design an algorithm to identify which defendants accepted a non-prosecution agreement. To do so, we check whether the sentence contains any expression connected to a non-prosecution agreement. This deterministic algorithm correctly classified almost all the cases into the category “defendant accepted a non-prosecution agreement”, making only five mistakes.
Finally, we design an algorithm to classify the remaining cases into two categories: “defendant is guilty” and “defendant was acquitted”. To do so, we define a bag of words that were found to be strong signals of acquittal and guilt when manually classifying the cases in our samples. [^15] Possebom (2023, Appendix I.2) lists the entire bag of words. We then count how many times each one of those expressions appears in each sentence, and we normalize those counts to be between 0 and 1.
Using the normalized counts, we train an L1-Regularized Logistic Regression using our training sample. We then validate this algorithm using our three validation samples and find that it correctly classifies 97.9% of the overall cases, 97.9% of the domestic violence cases, and 95.3% of the physical assault cases. Given this high success rate, we use the L1-Regularized Logistic Regression algorithm to define the treatment variable in our full sample.
Having designed the above algorithm, we use it to define the trial judge’s treatment variable $Y$ in the full sample. First, we find which defendants died during their trials and drop them from our sample. We then use the second and third algorithms to define which cases were dismissed and which cases are associated with a non-prosecution agreement. Moreover, we use the trained L1-regularized Logistic Regression algorithm to classify the remaining cases into the categories “defendant is guilty” and “defendant was acquitted”.
Appendix B Final-stage Judges Face Weaker Career Incentives
In Section 6, we find that only final-stage judges present a gender conviction rate gap for domestic violence cases and conclude that weaker career incentives for senior judges are one possible force driving the gender leniency gap. This interpretation relies on the assumption that early-stage judges face stronger career incentives than final-stage judges. The current appendix verifies the plausibility of this assumption.
To do so, we estimate the percentage of cases in which the assigned judge applied for a promotion opportunity within two years after the case assignment. Table B.1 shows these probabilities by crime type and judge characteristics. The first row focuses on the set of cases in which the assigned judge was not in the final stage of the Trial Court System, while the second row focuses on the set of cases in which the assigned judge was in the final stage of the Trial Court System. Odd columns restrict the sample to cases attributed to female judges while even columns restrict the sample to cases attributed to male judges. The header indicates the sample used to estimate these probabilities.
Table B.1: Percentage of Cases in which the Assigned Judge Applied for a Promotion Opportunity within Two Years by Crime Type and Judge Characteristics
Sample: Sample: Sample:
Judge Domestic Violence Physical Assault Misdemeanors
Career Stage Female Male Female Male Female Male
(1) (2) (3) (4) (5) (6)
Early 46.7% 50.9% 46.2% 49.2% 41.4% 48.4%
Final 0.0% 3.9% 0.0% 3.7% 0.0% 2.0%
  • Notes: This table shows the probability of each type of case being assigned to a judge who applied for a promotion opportunity within two years of the day the case was allocated to them. The first row shows this statistic for the set of cases that were attributed to judges who were not in the final stage of the Trial Court System (i.e., substitute judges or magistrates in the initial or intermediate stages). The second row shows this statistic for the set of crimes attributed to judges in the final stage of the Trial Court System. Odd columns restrict the sample to cases attributed to female judges. Even columns restrict the sample to cases attributed to male judges. Columns (1)-(2) refer to cases in the domestic violence sample. Columns (3)-(4) use the sample of physical assault cases. Columns (5)-(6) refer to cases from the misdemeanors sample.
We find that final-stage judges are much less likely to apply for promotion opportunities than early-stage judges. This result suggests that early-stage judges are more eager to be promoted and may face stronger career incentives than final-stage judges. Consequently, the assumption required to interpret weaker career incentives as one of the driving forces of the gender conviction rate gap is plausible.
Lastly, we briefly discuss whether the different conviction criteria adopted by male and female judges may have career consequences for the trial judges. As explained by Laneuville (2024), Appeals Court Judges vote on which trial judge to promote. Consequently, adopting sentencing criteria that lead to more appeal processes and reversed decisions may impact the judge’s career prospects.
To analyze the association between appeal processes, decision reversals, judges’ genders, and judges’ decisions, we regress an outcome variable indicating whether the trial judge’s decision was appealed or reversed on an indicator of a female judge, an indicator of a conviction decision, the interaction of the last two indicators, and district-by-quarter fixed effects. Table B.2 shows the results for the regression whose outcome variable indicates whether the trial judge’s decision was appealed while Table B.3 shows the results for the regression whose outcome variable indicates whether the trial judge’s decision was reversed.
Table B.2: Association between Conviction Sentences and Appealing of Judges’ Decisions on Domestic Violence Cases
Final Early
Appealed Appealed Appealed Appealed
(1) (2) (3) (4)
Fem $0.17^{***}$ $0.09^{***}$ $-0.01$ $0.03$
$(0.03)$ $(0.03)$ $(0.06)$ $(0.05)$
Convicted $0.45^{***}$ $0.41^{***}$
$(0.03)$ $(0.04)$
Fem x Convicted $-0.06$ $-0.09$
$(0.05)$ $(0.07)$
Mean men $0.21$ $0.21$ $0.29$ $0.29$
Quarter-District FE YES YES YES YES
Num. obs. $2,599$ $2,599$ $1,749$ $1,749$
${}^{***}p<0.01$; ${}^{**}p<0.05$; ${}^{*}p<0.1$
  • Notes: This table focuses on domestic violence cases and shows the association between a judge’s decision being appealed and the characteristics of the judges and of the decision. Columns (1) and (2) show the results for the sample of cases handled by final-stage judges. Columns (3) and (4) show the results for cases handled by early-stage judges.
Table B.2 shows that female senior judges face appeal processes more frequently than male judges (Column (1)) when analyzing domestic violence cases. It also shows that this higher probability of facing an appeal process is partially explained by convicting more cases (Column (2)). Moreover, it shows that early-stage judges, whose conviction probabilities do not vary by gender according to Table 7, present appeal probabilities that do not vary by gender either (Columns (3) and (4)). Table B.3 presents qualitatively similar results for the probability of having a decision reversed.
Table B.3: Association between Conviction Sentences and Reversal of Judges’ Decisions on Domestic Violence Cases
Final Early
Reversed Reversed Reversed Reversed
(1) (2) (3) (4)
Fem $0.02^{*}$ $0.02^{*}$ $-0.01$ $0.02$
$(0.01)$ $(0.01)$ $(0.06)$ $(0.02)$
Convicted $0.03^{***}$ $0.04^{**}$
$(0.01)$ $(0.01)$
Fem x Convicted $-0.01$ $-0.02$
$(0.02)$ $(0.02)$
Mean men $0.02$ $0.02$ $0.03$ $0.03$
Quarter-District FE YES YES YES YES
Num. obs. $2,599$ $2,599$ $1,749$ $1,749$
${}^{***}p<0.01$; ${}^{**}p<0.05$; ${}^{*}p<0.1$
  • Notes: This table focuses on domestic violence cases and shows the association between the reversal of a judge’s decision and the characteristics of the judges and of the decision. Columns (1) and (2) show the results for the sample of cases handled by final-stage judges. Columns (3) and (4) show the results for cases handled by early-stage judges.
We interpret these results as suggestive evidence that the gender conviction rate gap for domestic violence cases may have negative career consequences. For this reason, only female final-stage judges, who face weak career incentives, differ from their male counterparts when analyzing this type of case.

### v2

@@
 Abstract
-This article investigates the impact of judges’ gender on the outcome of domestic violence cases.
+We investigate the impact of judges’ gender on the outcome of domestic violence cases.
 Using data from São Paulo, Brazil, between 2010 and 2019, we compare conviction rates by judge’s gender and find that a domestic violence case assigned to a female judge is 31% (10 p.p) more likely to result in conviction than a case assigned to a male judge with similar career characteristics.
@@
 The United Nations Office on Drugs and Crime (2023) estimates that 55% of all homicides against women worldwide in 2022 were committed by intimate partners or other family members.
-As this violence is a criminal offense, a key component of the fight against domestic violence is the Criminal Court System and, therefore, how court judges approach this type of crime (Beecher-Monas, 2001).
+As this violence is a criminal offense, key components of the fight against domestic violence are the Criminal Court System and how court judges approach this type of crime (Beecher-Monas, 2001).
 For instance, the Human Rights Watch (1991) argues that, in the 1980s and early 1990s, Brazilian judges failed to analyze crimes of domestic violence against women in a non-discriminatory manner.
@@
 Physical assault cases are comparable to domestic violence cases because both involve some sort of physical aggression.
-We find that the gender conviction rate gap for domestic violence cases is more than twice the size of that of other misdemeanors.
-There are no differences between female and male judges in sentencing physical assault cases that do not involve domestic violence.
-Consequently, the gender conviction rate gap for domestic violence cases likely arises due to different perspectives about this type of crime, rather than because female judges are in general more punitive than their male counterparts.
-We find evidence of two drivers of the gender conviction rate gap.
-The first is sharing an identity with the typical victim of domestic violence cases.
-To investigate if the salience of gender identity matters, we test if the gender difference in strictness is more pronounced in intimate partner violence cases.
-The incidence and tolerance of this type of offense is strongly associated with patriarchal gender norms involving partners (Heise and Kotsadam, 2015; González and Rodríguez-Planas, 2020).
-To do this, we take advantage of two key features of our dataset.
-First, judges are randomly allocated to cases, which makes the relationship between the victim and the assailant independent of judge characteristics.
-Second, we are able to identify the victim–assailant relationship in cases that result in conviction.
-We find a statistically significant increase (37.5%)in the likelihood of conviction in a case involving intimate partner violence if the judge is a woman.
-However, the probability of conviction in a case not involving intimate partner violence is the same for men and women judges.
-Second, we test for gender differences in interpretation of the evidence provided in domestic violence cases.
-To test for the relevance of this driver, we compare the gender conviction rate between cases in which the defendant was caught by the police while committing the crime and cases without this objective piece of evidence.
-We find that male and female judges make similar decisions when analyzing cases where defendants are caught in the act, while judges differ when making decisions about cases with more subjective evidence.
-Hence, female judges interpret evidence differently from their male peers when analyzing domestic violence cases.
-This may mean that judges can make different decisions about what constitutes “true domestic violence” versus victims “gaming the system,” as characterized in Kafka et al.
-(2019), or that judges may disagree about the appropriate severity of punishment when there is room for discretion.
+We find that the gender conviction-rate gap for domestic violence cases is more than twice the size of that of other misdemeanors.
+Furthermore, while there is a gender conviction-rate gap for domestic violence cases, there are no differences between female and male judges when analyzing other types of physical assault cases.
+Consequently, the gender conviction-rate gap for domestic violence cases likely arises due to different perspectives about this type of crime rather than because female judges are, in general, more punitive than their male counterparts.
+For this reason, we classify this type of gap as an in-group bias according to the definition given by Shayo and Zussman (2011).
+We do so because victims of domestic violence are overwhelmingly female, and members of this group (female judges) act differently from members of the other group (male judges) when analyzing this type of crime specifically.
+We then investigate two possible drivers of this in-group bias.
+They were proposed by Boyd, Epstein, and Martin (2010) and are known as the representational account and the informational account.
+The first one argues that female judges act as representatives of their group and tend to protect it more intensely than male judges.
+The second one argues that female judges process the information contained in domestic violence cases in a unique way.
+Our evidence suggests that both accounts drive the in-group bias in domestic violence cases in São Paulo, Brazil.
+To investigate if the salience of gender identity matters (representational account), we test if the gender difference in strictness is more pronounced in intimate partner violence cases.
+To do so, we explore each case’s full sentence and estimate whether female judges are more likely to convict a defendant for a domestic violence offense and simultaneously flag the case as associated with intimate partner violence.
+We find that the gender conviction rate gap is 37.5% higher (9 p.p.) in cases in which a female judge may identify more strongly with the victim than a male judge, while there is no gender difference in cases that are not flagged as intimate partner violence offenses.
+This difference is associated with the representational account of the in-group bias because the incidence and tolerance of this type of offense are strongly associated with patriarchal gender norms involving partners (Heise and Kotsadam, 2015; González and Rodríguez-Planas, 2020).
+To test for gender differences in the interpretation of the evidence provided in domestic violence cases (informational account), we compare the gender conviction rate between cases in which the defendant was caught by the police while committing the crime and cases without this objective piece of evidence.
+We find that male and female judges make similar decisions when analyzing cases where defendants are caught in the act, while they differ when making decisions about cases with more subjective evidence.
+Hence, female judges process the information contained in domestic violence cases differently from their male peers, as suggested by the informational account of the in-group bias.
 Finally, we document that the difference in conviction rates between male and female judges is more pronounced later in their careers, which is when career pressures are weaker.
@@
 promotion is the main form of career incentive.
-However, promotion from the Trial Court to the Appeals Court is rare and extremely unlikely in the short run, unlike career progression at earlier stages.
-We find that there is no gender conviction rate gap for junior judges, while senior female judges convict domestic violence cases more frequently than male judges in the same career stage.
-This is an important avenue for future research because it could mean that the difference in perspective of early-career judges is not entirely reflected in the gender conviction gap, potentially because of early-career judges’ desire to conform to the court’s established norms.
-This article contributes to the literature about in-group bias in the judicial system.
-A large body of work finds evidence of judges making different decisions when they share a group identity with defendants or plaintiffs (Kruttschnitt and Savolainen, 2009; Shayo and Zussman, 2011a, 2017; Bielen and Grajzl, 2020; Cai et al., 2022; Chen et al., 2022).
-Similarly to our work, Kruttschnitt and Savolainen (2009), Bielen and Grajzl (2020), and Chen et al.
-(2022) analyze in-group gender bias in cases of uncommon crime types (rape and human trafficking), civil cases, and for female offenders, respectively.
+However, promotion from the Trial Court to the Appeals Court is rare and extremely unlikely in the short run, unlike career progression at earlier stages (Laneuville, 2024).
+We find that there is no gender conviction-rate gap for junior judges, while senior female judges convict domestic violence cases more frequently than male judges in the same career stage.
+This is an important avenue for future research because it could mean that the difference in perspective of early-career judges is not entirely reflected in the gender conviction-rate gap, potentially because of early-career judges’ desire to conform to the court’s established norms.
+This article contributes to the literature on judicial biases.
+Harris and Sen (2019) reviews this literature, focusing on ideology, race and gender gaps.
+Within the studies focusing on in-group bias, many authors find evidence of judges making different decisions when they share a group identity with defendants or plaintiffs (Kruttschnitt and Savolainen, 2009; Boyd, Epstein, and Martin, 2010; Shayo and Zussman, 2011, 2017; Bielen and Grajzl, 2020; Cai et al., 2022; Chen et al., 2022).
+Similarly to our work, Kruttschnitt and Savolainen (2009), Boyd, Epstein, and Martin (2010) Bielen and Grajzl (2020), and Chen et al.
+(2022) analyze in-group gender bias in cases of uncommon crime types (rape and human trafficking) in Finland, federal appellate cases in the U.S., civil cases in Belgium, and for female offenders in Kenya, respectively.
 Our work differs from theirs in that we focus on a single common crime type (domestic violence) whose victims are mostly women.
+Furthermore, while Boyd, Epstein, and Martin (2010) finds that in-group bias in federal appellate cases in the U.S.
+is consistent with the informational account only, we find that in-group bias in criminal cases in Brazil is consistent with the representational and informational accounts.
 The paper is organized as follows.
 Section 2 describes our dataset, while Section 3 explains our empirical strategy.
-In Section 4, we document the gender conviction rate gap, while 5 discusses its driving forces.
-Section 7 concludes.
+Moreover, Section 4 documents the gender conviction rate gap, while Section 5 discusses its driving forces.
+Finally, Section 6 argues that the difference in perspective of early-career judges may not be entirely reflected in the gender conviction-rate gap, and Section 7 concludes.
 Data Description and Summary Statistics
@@
 The data also contain two pieces of information that are crucial for the construction of our sample, which we discuss in detail in Section 2.3.
-The first is a dummy indicating if the case was randomly assigned to a judge within a court district and point in time [^3] That is, not connected to other crime randomly assigned previously..
-The second is the name of the courthouse, where different courthouses correspond to certain legal specializations.
+The first is a dummy indicating if the case was randomly assigned to a judge within a court district and point in time.
+[^3] Alternatively, some cases are connected to other crimes and are assigned to the judge who was assigned to the previous criminal case.
+The second is the name of the courtroom, where different courtrooms may correspond to certain legal specializations.
 We use this information to make sure that domestic violence cases are indeed randomly assigned within a district and not systematically assigned to (or reassigned to) judges who do not preside over cases of other types of crimes.
@@
 First, we restrict the dataset to cases that were randomly allocated to judges.
-By Brazilian Law, a case can be allocated to judges by either connected distribution and free distribution.
+By Brazilian Law, a case can be allocated to judges by either connected distribution or free distribution.
 In the first procedure, the crime under consideration is connected to an existing criminal process.
 Here, the current case is then attached to the previous process and analyzed by the judge associated with the older case.
-[^8] For two simple examples, a case where a murderer then stole a car to flee the police or a case where a drug trafficker was then caught laundering money.
+[^8] For two simple examples, consider a case where a murderer killed someone and then stole a car to flee the police or a case where a drug trafficker was also caught laundering money.
 In the second procedure, the crime under consideration is not connected to an existing criminal process.
@@
 Last, we need to ensure that the assigned judge’s gender is determined correctly.
-To do this, we focus on cases were allocated to judges between January 2011 and December 2019 and use our productivity dataset to remove all observations in which the assigned judge was not active at the court district and month of the assignment.
+To do this, we focus on cases allocated to judges between January 2011 and December 2019 and use our productivity dataset to remove all observations in which the assigned judge was not active at the court district and month of the assignment.
 [^9] Cases are randomly allocated to courtrooms and instead of directly to individual judges.
-If a court district is understaffed at the time of case assignment and there is no presiding judge in the assigned courtroom assigned, the system may record the case as being attributed to the next person who takes over the courtroom, which may take place at a later point in time.
+If a court district is understaffed at the time of case assignment and there is no presiding judge in the assigned courtroom, the system may record the case as being attributed to the next person who takes over the courtroom, which may take place at a later point in time.
 Using this final sample, we compute summary statistics for each possible case outcome and crime type.
-Table 1, displays case outcomes in the columns and crime types in the rows.
+Table 1 displays case outcomes in the columns and crime types in the rows.
+It also reports the share of female judges analyzing each type of crime.
 Domestic violence cases are more frequently acquitted than other misdemeanors or physical assault cases.
 As conviction rates are similar across types of crimes, this higher acquittal is due to fewer NPAs and dismissed cases.
-Table 1: Case Outcomes by Crime Type in %
-Crime Type Conviction NPA Acquittal Dismissal Expired Dead # Cases
-(1) (2) (3) (4) (5) (6)
-DV 32.5 1.1 51.4 4.3 8.8 2.0 4,358
-Other Mis 36.9 5.4 31.2 15.5 9.5 2.2 39,815
-Other PA 27.2 7.0 34.1 20.6 10.4 1.9 1,648
+We also note that the share of female judges is similar across crime types and that there are 270 female judges and 425 male judges in our entire sample.
+Table 1: Case Outcomes and Share of Female Judges by Crime Type in %
+Crime Conviction NPA Acquittal Dismissal Expired Dead Female Obs
+Type Judge
+DV 32.5 1.1 51.4 4.3 8.8 2.0 32.1 4,358
+Other Mis 36.9 5.4 31.2 15.5 9.5 2.2 36.4 39,815
+Other PA 27.2 7.0 34.1 20.6 10.4 1.9 38.1 1,648
   • Notes: The rows of this table indicate the different samples used in this paper: domestic violence cases (DV), misdemeanor cases unrelated to domestic violence (Mis), and physical assault cases that are not domestic violence (PA).
@@
 The “expired” dummy equals one if at least one of the charges in the case has expired.
-The “Dead” dummy equals one if the defendant died the sentence’s date.
+The “Dead” dummy equals one if the defendant died before the sentence’s date.
 The rows do not sum to 100% because cases with multiple charges can have different outcomes for each charge.
+The second to last column shows the share of female judges in each sample while the last column shows the size of each sample.
 Empirical Strategy and Validity Tests
-In this section, we explain the empirical strategy used to measure the gender conviction rate gap and indirectly test the strategy’s validity.
+In this section, we explain the empirical strategy used to measure the gender conviction-rate gap and indirectly test the strategy’s validity.
 We measure the gender conviction rate gap for domestic violence cases by comparing the conviction probability for similar cases allocated to female and male judges.
@@
 Our parameter of interest, $\beta$, measures whether a female judge is more likely to convict than a male judge under the assumption that the judge’s gender is independent of the case’s potential outcome.
-This exogeneity assumption is valid according to Brazilian Law mandating cases are randomly allocated to judges within court districts (Section 2.3).
+This exogeneity assumption is valid according to Brazilian Law, which mandates cases to be randomly allocated to judges within court districts (Section 2.3).
 We indirectly test the validity of this assumption in Table 2.
 Even when our exogeneity assumption is valid, the $\beta$ parameter in Equation (1) may capture differences between male and female judges that are unrelated to their views on domestic violence.
-To measure whether this conviction rate gap is due to different gender perspectives about this type of crime, we compare this domestic violence conviction gap against the gap for similar crime types.
-Our sample therefore consists of domestic violence cases and one type of comparable criminal cases.
+To measure whether this conviction-rate gap is due to different gender perspectives about this type of crime, we compare this domestic violence conviction gap against the gap for similar crime types.
+Now, our sample consists of domestic violence cases and one type of comparable criminal case.
 We estimate the following linear regression model:
@@
 In Equation (2), our parameter of interest is $\beta_{3}$.
-This specification accounts for the extra strictness and social unacceptability of domestic violence offenses beyond that of other violent crimes of comparable severity.
-Consequently, this coefficient captures the part of the gender conviction rate gap that is likely due to different gender perspectives about domestic violence cases.
+This specification accounts for the extra strictness and social unacceptability of domestic violence offenses beyond that of crimes of comparable severity (coefficient $\beta_{2}$) and for the possibility that female judges are stricter than male judges in general (coefficient $\beta_{1}$).
+Consequently, this coefficient captures the part of the gender conviction-rate gap that is likely due to different gender perspectives about domestic violence cases specifically.
+For this reason, we interpret the $\beta_{3}$ coefficient as capturing a type of in-group bias according to the definition given by Shayo and Zussman (2011).
+We do so because victims of domestic violence are overwhelmingly female, and members of this group (female judges) may act differently from members of the other group (male judges) when analyzing this type of crime in particular.
 This interpretation for the $\beta_{3}$ coefficient is valid if the judge’s gender is independent of the case’s potential outcome after conditioning on the type of crime.
@@
 Sample Physical Assault Misdemeanors Domestic Violence
-Quarter x Court District FE YES YES YES
+Quarter x Court District FE $\checkmark$ $\checkmark$ $\checkmark$
 Num. obs. $6,006$ $44,173$ $4,358$
@@
   • Notes: Each cell represents the linear relationship between the gender of the judge randomly assigned to the case and a set of characteristics of the case.
-Column 1 shows this relationship for the sample of all physical assault crimes, Column 2 shows this relationship for the sample of misdemeanors (crimes such that the maximum sentence does not exceed 4 years).
-Column 3 shows this relationship for the sample of domestic violence crimes.
+Column (1) shows this relationship for the sample of all physical assault crimes, Column (2) shows this relationship for the sample of misdemeanors (crimes such that the maximum sentence does not exceed 4 years), and Column (3) shows this relationship for the sample of domestic violence crimes.
 All regressions include quarter-by-court district fixed effects.
-Measuring the Gender Conviction Rate Gap
+Measuring the Gender Conviction-Rate Gap
 We measure the gender conviction rate gap using two empirical strategies.
@@
 Table 3 shows the effect of judge gender on the probability of conviction in domestic violence cases according to Equation (1).
-In the first column, we find that randomly selecting a woman or a man to preside does significantly impact conviction rates.
+In the first column, we find that randomly selecting a woman or a man to preside does not significantly impact conviction rates.
 However, as shown in Laneuville (2024), the average male judge and the average female judge have different career trajectories.
@@
 $(0.03)$ $(0.03)$ $(0.03)$
-Experience $0.01^{***}$ $0.01^{***}$
-$(0.00)$ $(0.00)$
-Years in Career Stage $-0.01^{*}$ $-0.01^{*}$
-$(0.01)$ $(0.01)$
-Caught Red Handed? $0.06$
-$(0.04)$
-Defendant in jail $-0.03$
-$(0.04)$
 Male Judge’s Conviction Rate $0.32$ $0.32$ $0.32$
-Quarter x Court District FE YES YES YES
-Career Stage Controls NO YES YES
+Quarter x Court District FE $\checkmark$ $\checkmark$ $\checkmark$
+Career Controls $\checkmark$ $\checkmark$
+Case Controls $\checkmark$
 Num. obs. $4,358$ $4,358$ $4,358$
@@
 The second column includes controls for judge characteristics (career stage, number of years of experience as a judge, and number of years of experience in the current career stage).
-The third column controls for both judge and case characteristics.
+The third column controls for both judge and case characteristics (whether the defendant was caught red-handed and whether the defendant was arrested by the sheriff before the trial).
 All regressions include quarter-by-court district fixed effects.
@@
 Men’s CR (All) $0.37$ $0.37$ $0.37$ $0.31$ $0.31$ $0.31$
-Career Controls NO YES YES NO YES YES
-Case Controls NO NO YES NO NO YES
-Quarter-District FE YES YES YES YES YES YES
+Quarter-District FE $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+Career Controls $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+Case Controls $\checkmark$ $\checkmark$
 Num. obs. $44,173$ $44,173$ $44,173$ $6,006$ $6,006$ $6,006$
@@
 The interaction coefficient captures the part of the gender conviction rate gap that is specific to domestic violence cases and is likely due to different approaches to domestic violence cases between male and female judges.
-In the next section, we explore this difference more deeply.
+For this reason, this result suggests the existence of a type of in-group bias according to the definition given by Shayo and Zussman (2011).
+[^10] In Appendix C, we propose a simple economic model formalizing the definition of in-group bias.
+In the next section, we explore the driving forces behind this in-group bias.
 Understanding the Gender Conviction Rate Gap
-In this section, we show two important findings about the conviction rate gender gap.
-The first is suggestive evidence that the gap is explained by judges’ identification with their own gender.
-We find that the difference in conviction rates between male and female judges is more intense in domestic violence cases involving intimate partner violence.
-Violence between relationship partners is highly impacted by patriarchal social norms that make gender identity more salient.
+In this section, we show two important findings about the conviction rate gender gap and its driving forces.
+The first is suggestive evidence that the gap is explained by judges’ identification with their own gender (Section 5.1).
+We find that female judges are more likely than male judges to convict and simultaneously flag the case as an intimate partner violence offense, while female and male judges are equally likely to convict and not flag the case as an intimate partner violence offense.
+This result suggests that the representational account (Boyd, Epstein, and Martin, 2010) is one of the explanatory forces of the in-group bias documented in Table 4 because violence between relationship partners is highly impacted by patriarchal social norms that make gender identity more salient (Heise and Kotsadam, 2015; González and Rodríguez-Planas, 2020).
 The second finding, discussed in Section 5.2, is that men and women respond differently to evidence presented in the same case.
 We present evidence that female judges are more likely than male judges to convict in cases where evidence is more tenuous and subject to interpretation, but not when the evidence is objectively incriminating.
+This result suggests that the informational account (Boyd, Epstein, and Martin, 2010) is one of the explanatory forces of the in-group bias documented in Table 4.
+In Appendix C, we propose a simple threshold-crossing model that formalizes the concepts of the representational and informational accounts.
 Female Identity as a Driver for Conviction Rates
-Next, we show that the gender of the judge matters more in cases where patriarchal social norms are particularly relevant: intimate partner violence.
-We interpret this finding as indicative that a major reason for the gender conviction gap is that domestic violence cases make gender identity more salient to judges.
-First, sexist social norms and gender roles significantly affect the occurrence and punishment of intimate partner violence in multiple contexts.
+In this section, we show that the gender of the judge matters more in cases where patriarchal social norms are particularly relevant: intimate partner violence.
+We interpret this finding as indicative that a major reason for the gender conviction gap is that domestic violence cases make gender identity more salient to judges, suggesting that the representational account (Boyd, Epstein, and Martin, 2010) is one of the explanatory forces of the in-group bias documented in Section 4.
+We argue that domestic violence cases make gender identity more salient because sexist social norms and gender roles significantly affect the occurrence and punishment of intimate partner violence in multiple contexts.
 For instance, González and Rodríguez-Planas (2020) show that, among immigrant families, the incidence of intimate partner violence is positively affected by gender norms in the country of origin.
@@
 Closer to our context, Perova and Reynolds (2017) show that police stations that specialize in domestic violence with specially trained officers significantly reduce the number of deaths of young women by intimate partner violence in metropolitan municipalities in Brazil.
-In terms of judicial decision-making on domestic violence, the negative stereotypes referred to in Kafka et al.
+In terms of judicial decision-making on domestic violence, the negative stereotypes mentioned by Kafka et al.
 (2019) and Human Rights Watch (1991) concern female intimate partners.
-We then take advantage of an important feature of our data: we are able to identify the relationship between the victim and the assailant for all cases that culminate in conviction.
-[^10] In fact, we observe this relationship for cases in which the judge must make an active decision, e.g., excluding cases where the judge declares the defendant dead, a non-prosecution agreement is reached, or the case expires.
+To show that the gender of the judge matters more for the outcome of intimate partner violence cases than for that of other domestic violence offenses, we take advantage of an important feature of our data.
+We are able to identify the relationship between the victim and the assailant for all cases that culminate in an active decision by the presiding judge (conviction or acquittal).
 To do this, we construct an algorithm that defines a case as intimate partner violence if, when writing the court’s decision, the judge uses the Portuguese word for wife, girlfriend, ex-wife, ex-girlfriend, or similar words.
@@
 The specific words we look for are “namorada,” “ex-namorada,” “esposa,” “ex-esposa,” “amásia,” “ex-amásia,” “companheira,” “ex-companheira,” “mulher,”, “ex-mulher,” “coabitação,” “união estável,” “casamento,” “era casado,” and “era casada.”
-To show that the gender of the judge matters more for the outcome of intimate partner violence cases than for that of other domestic violence offenses, we make the following argument.
-If the probability of a conviction for non-intimate partner violence is non-zero, then two things cannot be true at the same time: 1) the probability of conviction for a domestic violence offense involving intimate partner violence is higher for women than it is for men; 2) the probability of conviction for a domestic violence offense not involving an intimate partner is not higher for women than it is for men.
-Although we cannot observe the relationship between victim and assailant in all cases, we can reasonably assume that this relationship is independent of the assigned judge’s gender due to the fact that cases are randomly allocated to judges in our sample.
-To test for this possibility, we interact the indicator for assault between intimate partners with the conviction indicator. This interaction term becomes our new outcome variable in Equation (1).
+We, then, test whether female judges are more likely to make a conviction decision in a domestic violence case and flag the case as being related to intimate partner violence after controlling for case characteristics.
+If there is a gender conviction gap in cases flagged as intimate partner violence but not in cases without such a flag, then a stronger identification with the victim may be driving the gender lenience gap found in Section 4.
+If the gender conviction rate gap is present regardless of explicit mentions of intimate partner violence, then there must be other forces driving this lenience gap.
+To implement this test, we interact the indicator for assault between intimate partners with the conviction indicator. This interaction term becomes our new outcome variable in Equation (1).
 Table 5 shows the results of these regressions.
@@
 Men’s Outcome Average $0.24$ $0.07$ $0.32$
-Quarter-District FE YES YES YES
-Case Controls YES YES YES
-Judge Career Controls YES YES YES
+Quarter-District FE $\checkmark$ $\checkmark$ $\checkmark$
+Case Controls $\checkmark$ $\checkmark$ $\checkmark$
+Judge Career Controls $\checkmark$ $\checkmark$ $\checkmark$
 Num. obs. $4,358$ $4,358$ $4,358$
@@
 All regressions include quarter-district fixed effects and a full set of controls for case characteristics and judge career stages.
-These findings are consistent with existing evidence that, in many contexts, judges make different decisions when they share a group identity with defendants or plaintiffs (Kruttschnitt and Savolainen, 2009; Shayo and Zussman, 2011a, 2017; Bielen and Grajzl, 2020; Cai et al., 2022; Chen et al., 2022).
+These findings are consistent with existing evidence that, in many contexts, judges make different decisions when they share a group identity with defendants or plaintiffs (Kruttschnitt and Savolainen, 2009; Shayo and Zussman, 2011, 2017; Bielen and Grajzl, 2020; Cai et al., 2022; Chen et al., 2022).
 Many of these papers show that the differences intensify as group identity becomes more salient.
 This intensification may result from of a heightened opposition between two or more social groups (“us” versus “them”).
-For example, Shayo and Zussman (2011b) show that Jewish Israeli judges in small claim courts favor Jewish litigants more intensely when they are more exposed to conflict-related terrorism.
+For example, Shayo and Zussman (2011) show that Jewish Israeli judges in small claim courts favor Jewish litigants more intensely when they are more exposed to conflict-related terrorism.
 Closer to our case, Cai et al.
 (2022) show that judges analyzing divorce cases in China favor plaintiffs of their own gender more in areas where social norms are more conservative.
-In our case, we show that the gender gap is much more intense in cases where the form of domestic violence is more stereotypical [^12] Note that most domestic violence against women is committed by partners and ex-partners (United Nations Office on Drugs and Crime, 2023).
-and has the potential to cause the most division because of social norms.
-Judges’ views on other forms of domestic violence may be less dependent on their views on patriarchy.
-For example, beating one’s own parents or children is more likely to be viewed as unjustifiable by men and women alike.
+In our case, we show that the gender gap is much more intense in cases where the form of domestic violence is more stereotypical.
+[^12] Note that most domestic violence cases against women are committed by partners and ex-partners (United Nations Office on Drugs and Crime, 2023).
+Consequently, the representational account (Boyd, Epstein, and Martin, 2010) is a plausible driving forcing of the in-group bias documented in Section 4.
 Male and Female Judges Analyze Evidence Differently
+In this section, we show that female judges are more likely than male judges to convict in domestic violence cases where evidence is more tenuous and subject to interpretation, but not when the evidence is objectively incriminating.
+We interpret this result as suggesting that the informational account (Boyd, Epstein, and Martin, 2010) is one of the explanatory forces of the in-group bias documented in Table 4.
 Domestic violence cases require complex evidentiary rulings.
@@
 These cases have more objective evidence, often including testimony from police officers who witnessed the offense.
-If the gender conviction rate gap is stronger in cases where the defendant was not caught red-handed, then this may be because female and male judges interpret more subjective pieces of evidence differently.
-This would provide possible explanatory force for the observed leniency gap in domestic violence cases.
+If the gender conviction rate gap is stronger in cases where the defendant was not caught red-handed, then this may be because female and male judges interpret more subjective pieces of evidence differently (informational account).
 If the gender conviction rate gap is similar across domestic violence cases regardless of the type of evidence presented, then there must be other forces driving this lenience gap.
@@
 Men’s CR (All) $0.35$ $0.30$ $0.47$ $0.32$ $0.36$ $0.30$
-Quarter-District FE YES YES YES YES YES YES
-Judge Career Controls YES YES YES YES YES YES
+Quarter-District FE $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+Judge Career Controls $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 Num. obs. $1,337$ $3,021$ $14,391$ $29,782$ $1,552$ $4,454$
@@
 These results suggest that female and male judges interpret less objective evidence differently, while they interpret objective evidence (cases flagged as in flagrante delicto) in similar ways.
-These different interpretations of more subjective cases offer one explanation for the gender conviction rate gap in domestic violence cases.
+These different interpretations of more subjective cases are consistent with the informational account being one possible driving force of the in-group bias documented in Section 4.
 Discussion: Career Incentives and a Potential Underestimation of the Gap
-Next, we analyze the association between the strength of career incentives and the size of the conviction gender gap.
+In this section, we analyze the association between the strength of career incentives and the size of the conviction gender gap.
 We show that male and female judges differ more intensely at the end of their careers.
 This corresponds with the period where they face the lowest probability of upcoming performance evaluation for promotion to the Appeals Court.
-[^13] See Appendix B for more information on the proportion of cases analyzed by judges who apply for promotion in the near future Career pressures may lead early-career male and female judges to act more similarly, possibly in response to feedback from Appeals Court magistrates, who are overwhelmingly male.
+[^13] See Appendix B for more information on the proportion of cases analyzed by judges who apply for promotion in the near future Career pressures may lead early-career male and female judges to act more similarly, possibly in response to feedback from Appeals Court magistrates, who are overwhelmingly male (Laneuville, 2024).
 Since Brazilian judges cannot be fired and must receive the same salary, promotion is the main career incentive faced by magistrates in our sample.
@@
 Men’s CR (All) $0.26$ $0.43$ $0.35$ $0.50$ $0.25$ $0.50$
-Quarter-District FE YES YES YES YES YES YES
-Case Controls YES YES YES YES YES YES
+Quarter-District FE $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+Case Controls $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 Num. obs. $2,609$ $1,749$ $25,639$ $18,534$ $3,499$ $2,507$
@@
 Conclusion
-In this paper, we provide evidence that cases assigned to female judges are significantly more likely to result in conviction than those assigned to a male judges with a similar career path.
+In this paper, we provide evidence that cases assigned to female judges are significantly more likely to result in conviction than those assigned to male judges with similar career paths.
 This pattern is, to a large extent, due to the reasons that are specific to domestic violence.
-The gender difference in conviction rates is stronger for cases involving domestic violence than for crimes that are similar in nature (physical assault) and sentence length (misdemeanors).
+The gender difference in conviction rates is stronger for cases involving domestic violence than for crimes that are similar in nature (physical assault) and sentence length (misdemeanors), suggesting the existence of a type of in-group bias as defined by Shayo and Zussman (2011).
 Furthermore, we find suggestive evidence for two major drivers of the conviction gender gap in domestic violence.
@@
 We claim this is evidence of identification with one’s own gender, as perception of intimate partner violence is highly affected by patriarchal social norms, as opposed to other forms of violence where this might not be the case.
+This result is consistent with the representational account as defined by Boyd, Epstein, and Martin (2010).
 The second driver comes from gender differences in the way judges analyze evidence.
 There is no gender gap when the defendant is caught red-handed by the police, making the evidence presented clear and objective, but the gap is very strong when there is more room for subjectivity in the analysis.
+This result is consistent with the informational account as defined by Boyd, Epstein, and Martin (2010).
 Finally, we identify an association between a larger gender gap and weaker career pressures for more senior judges in the final stage of the trial court.
@@
 Mean men $0.21$ $0.21$ $0.29$ $0.29$
-Quarter-District FE YES YES YES YES
+Quarter-District FE $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 Num. obs. $2,599$ $2,599$ $1,749$ $1,749$
@@
 Mean men $0.02$ $0.02$ $0.03$ $0.03$
-Quarter-District FE YES YES YES YES
+Quarter-District FE $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 Num. obs. $2,599$ $2,599$ $1,749$ $1,749$
@@
 For this reason, only female final-stage judges, who face weak career incentives, differ from their male counterparts when analyzing this type of case.
+Appendix C Modelling the Gender Conviction-Rate Gap
+In this appendix, we propose a simple economic model to formalize the concepts of the gender conviction-rate gap, in-group bias (Shayo and Zussman, 2011) and their driving forces, the representational and informational accounts (Boyd, Epstein, and Martin, 2010).
+First, to formally define the first two concepts, we adapt the potential outcome model (Rubin, 1974) to a scenario where cases of different types of crime may be treated with a female judge instead of a male judge.
+Second, to formally define the representational and informational accounts of the in-group bias, we adapt the threshold-crossing model (Heckman and Vytlacil, 2005) to a context where the latent heterogeneity is associated with the evidence of a criminal case and the propensity score is associated with the judge’s conviction criteria.
+We start by looking at criminal cases as our observational unit.
+Each case is associated with a type of crime: $d=1$ indicates that it is a domestic violence offense, and $d=0$ indicates that it is either a misdemeanor offense or another type of physical assault case.
+Moreover, each case may receive two randomly allocated treatments: $f=1$ indicates that the case was assigned to a female judge and $f=0$ indicates that the case was assigned to a male judge.
+Finally, each case has a potential conviction decision, $Y\left(f,d\right)$, which depends on the gender of the presiding judge and on the type of crime.
+[^16] Note that, in our context, sex is a manipulable variable because the unit of observation is the case instead of the judge.
+Moreover, type of crime is a pre-determined covariate, because it is chosen by the prosecutor before the presiding judge is assigned.
+We define the gender conviction-rate gap as
+$\mathbb{E}\left[Y\left(1,1\right)-Y\left(0,1\right)\right]\neq 0,$ (C.1)
+i.e., the probability that a domestic violence case ends in a conviction when presided by a female judge is different from the probability that a domestic violence case ends in a conviction when presided by a male judge.
+We also define the in-group bias as
+$\left\{\begin{matrix}[l]\mathbb{E}\left[Y\left(1,1\right)-Y\left(0,1\right)\right]&>&0\\ \mathbb{E}\left[Y\left(1,1\right)-Y\left(0,1\right)\right]&>&\mathbb{E}\left[Y\left(1,0\right)-Y\left(0,0\right)\right]\end{matrix}\right.,$ (C.2)
+i.e., the gender conviction-rate gap for domestic violence cases (i) is positive, and (ii) is larger than the gender conviction-rate gap for other types of crime.
+Next, we formally define the representational and informational account using our empirical context and a threshold-crossing model.
+In any criminal case, evidence is produced against the defendant.
+In our model, the judge sets a conviction criterion and interprets the evidence.
+Then, the judge convicts the defendant if the judge interprets the evidence as being sufficient to do so.
+The evidence interpretation function, $E\colon\left\{0,1\right\}^{3}\times\mathbb{R}\rightarrow\mathbb{R}$, depends on the judge’s gender (a female indicator variable, $F$), the type of crime (a domestic violence indicator variable, $D$), whether there is objective and strong evidence against the defendant (an indicator variable $R$ that is equal to one if the defendant was caught red-handed when committing the crime), and all other types of evidence produced by the police and the prosecution, $V$.
+Hence, $E\left(F,D,R,V\right)$ is a random variable capturing the amount of evidence against the defendant as interpreted by the judge assigned to the case.
+The judge’s conviction criterion function, $P\colon\left\{0,1\right\}^{3}\rightarrow\mathbb{R}$, depends on the judge’s gender, the type of crime and whether the crime is associated with intimate partner violence (an indicator variable $W$ that is equal to one if the victim is the wife or intimate partner of the defendant).
+Hence, $P\left(F,D,W\right)$ is a random variable capturing the conviction criterion of the judged assigned to the case.
+The judge convicts the defendant if the judge interprets the evidence as being sufficient to do so, i.e.,
+$Y=\mathbf{1}\left\{P\left(F,D,W\right)\leq E\left(F,D,R,V\right)\right\},$ (C.3)
+where $Y$ denotes whether the defendant is convicted or not.
+Note that Equation (C.3) imposes two exclusion restrictions.
+While the relationship status between the victim and the defendant — $W$ —- only enters the judge’s conviction criterion function, being caught red-handed — $R$ — only enters the evidence interpretation function.
+The first exclusion restriction is imposed because the Criminal Process Code does not include this variable as valid evidence against the defendant, but the Criminal Law Code includes this variable as an aggravating factor when setting the sentence.
+The second exclusion restriction is imposed because the Criminal Process Code considers in flagrante delicto as a piece of evidence against the defendant but not as an aggravating factor.
+Using this model, we can formally define the representational account driving the in-group bias as
+$P\left(1,1,1\right)-P\left(0,1,1\right)<P\left(1,1,0\right)-P\left(0,1,0\right)\leq 0,$ (C.4)
+i.e., female judges may set stricter conviction criteria than male judges for domestic violence cases and female judges are even stricter when the case is associated with intimate partner violence.
+We can also formally define the informational account driving the in-group bias as
+$E\left(1,1,0,v\right)-E\left(0,1,0,v\right)>E\left(1,1,1,v\right)-E\left(0,1,1,v\right)\geq 0,$ (C.5)
+for any $v$ in the support of $V$.
+When these inequalities hold, the female judge interprets the evidence more harshly than a male judge when the defendant is caught red-handed and the female judge interprets the evidence even more harshly when the defendant is not caught red-handed.
+In the main text, we use data from criminal offenses in the State of Sao Paulo, Brazil, to test the existence of each one of the four concepts defined above.
+In Table 3, we show that the gender conviction-rate gap (Equation (C.1)) exists.
+Then, in Table 4, we show that in-group bias (Equation (C.2)) is present in our application.
+Finally, Table 5 shows that our data is consistent with the representational account (Equation (C.4) while Table 6 provides evidence that our data is consistent with the informational account (Equation (C.5).

### v3

@@
 We investigate the impact of judges’ gender on the outcome of domestic violence cases.
-Using data from São Paulo, Brazil, between 2010 and 2019, we compare conviction rates by judge’s gender and find that a domestic violence case assigned to a female judge is 31% (10 p.p) more likely to result in conviction than a case assigned to a male judge with similar career characteristics.
-To show that this decision gap rises due to different gender perspectives about domestic violence instead of rising due to female judges being tougher than male judges, we compare it against gender conviction rate gaps in similar types of crimes.
-We find that the gender conviction rate gap for domestic violence cases is significantly larger than the same gap for other misdemeanor cases (3 p.p.
+Using data from São Paulo, Brazil, between 2010 and 2019, we compare conviction rates by judge’s gender and find that a domestic violence case assigned to a female judge is 31% (10 p.p.) more likely to result in conviction than a case assigned to a male judge with similar career characteristics.
+To show that this decision gap rises due to different gender perspectives about domestic violence and not because female judges are stricter than their male counterparts in all rulings, we compare it against the gender conviction-rate gap in similar types of crime.
+We find that the gender conviction-rate gap for domestic violence cases is significantly larger than the same gap for other misdemeanor cases (3 p.p.
 larger) and for other physical assault cases (8 p.p.
 larger).
-Lastly, we find evidence that at least two channels explain this gender conviction rate gap for domestic violence cases: gender-based differences in evidence interpretation and gender-based sentencing criteria.
+Furthermore, we find evidence that at least two channels explain this gender conviction-rate gap for domestic violence cases: gender-based differences in evidence interpretation and gender-based sentencing criteria.
+Lastly, we find that this gender conviction rate has no significant impact on the probability of appeals, ruling reversals or recidivism.
 Keywords: Domestic Violence, Judicial Decisions, Gender Bias
@@
 The most prominent of these is the Maria da Penha Law, enacted in 2006, which characterizes domestic violence as a specific type of criminal offense.
+Ferraz and Schiavon (2022) find that this law reduced female homicides due to household aggression by 9 percent, with stronger effects for less educated and black women.
 Moreover, Kafka et al.
@@
 Since discriminatory gender stereotypes may vary depending on the decision maker’s individual characteristics, we investigate the impact of judges’ subjectivity on the outcome of domestic violence cases.
-In particular, we focus on the effect of a judge’s gender on their decisions, documenting a gender conviction-rate gap specific to domestic violence cases.
-We also analyze possible forces driving this difference.
-To do this, we collect data from São Paulo, Brazil, between 2010 and 2019 and compare conviction rates by the deciding judge’s gender after controlling for the judge’s career characteristics.
+In particular, we focus on the effect of a judge’s gender on their decisions, documenting a sizeable and significant gender conviction-rate gap specific to domestic violence cases.
+We also analyze two possible forces driving this difference.
+At the end, we discuss the consequences of this gender conviction-rate gap beyond the Trial Court, analyzing the outcome of domestic violence cases in the Appeals Court and the recidivism behavior of these cases’ defendants.
+We find that the Trial Judge’s gender has no effect on these outcomes.
+To do this, we collect data from São Paulo, Brazil, between 2010 and 2019 and compare conviction rates by the presiding judge’s gender after controlling for the judge’s career characteristics.
 Case characteristics are independent of the presiding judge’s gender because cases are randomly allocated to judges within each court district in Brazil.
@@
 In this analysis, we find that a domestic violence case assigned to a female judge is 31% (10 p.p) more likely to result in conviction than a case assigned to a male judge with similar career characteristics.
-Since this decision gap may be due to differences between male and female judges regardless of their perspectives on domestic violence, we compare this gap to gender decision gaps for two other types of crime: misdemeanors and physical assault.
+Since this decision gap may be due to differences between male and female judges regardless of their perspectives on domestic violence, we compare this gap to the gender decision gaps for two other types of crime: misdemeanors and physical assault cases.
 Because we focus on domestic violence crimes that do not result in death or long-term physical impairment, misdemeanors are comparable to domestic violence cases in their sentence severity (according to Brazilian Law).
@@
 We find that the gender conviction-rate gap for domestic violence cases is more than twice the size of that of other misdemeanors.
-Furthermore, while there is a gender conviction-rate gap for domestic violence cases, there are no differences between female and male judges when analyzing other types of physical assault cases.
+Furthermore, while there is a gender conviction-rate gap for domestic violence cases, there are no differences between female and male judges when analyzing other types of physical assault.
 Consequently, the gender conviction-rate gap for domestic violence cases likely arises due to different perspectives about this type of crime rather than because female judges are, in general, more punitive than their male counterparts.
@@
 To do so, we explore each case’s full sentence and estimate whether female judges are more likely to convict a defendant for a domestic violence offense and simultaneously flag the case as associated with intimate partner violence.
-We find that the gender conviction rate gap is 37.5% higher (9 p.p.) in cases in which a female judge may identify more strongly with the victim than a male judge, while there is no gender difference in cases that are not flagged as intimate partner violence offenses.
-This difference is associated with the representational account of the in-group bias because the incidence and tolerance of this type of offense are strongly associated with patriarchal gender norms involving partners (Heise and Kotsadam, 2015; González and Rodríguez-Planas, 2020).
+We find that the gender conviction-rate gap is 37.5% higher (9 p.p.) in cases in which a female judge may identify more strongly with the victim than a male judge and statistically null in cases that are not flagged as intimate partner violence offenses.
+This difference is associated with the representational account of the in-group bias because the incidence and tolerance of this type of offense are strongly associated with patriarchal gender norms involving partners (Heise and Kotsadam, 2015; Tur-Prats, 2019; González and Rodríguez-Planas, 2020).
 To test for gender differences in the interpretation of the evidence provided in domestic violence cases (informational account), we compare the gender conviction rate between cases in which the defendant was caught by the police while committing the crime and cases without this objective piece of evidence.
-We find that male and female judges make similar decisions when analyzing cases where defendants are caught in the act, while they differ when making decisions about cases with more subjective evidence.
+We find that male and female judges make similar decisions when analyzing cases where defendants are caught in the act, while they differ when making decisions about cases that rely on more subjective evidence.
 Hence, female judges process the information contained in domestic violence cases differently from their male peers, as suggested by the informational account of the in-group bias.
-Finally, we document that the difference in conviction rates between male and female judges is more pronounced later in their careers, which is when career pressures are weaker.
-We define late-career judges as those in the last stage of the career of a Trial Court judge.
-Because judges in Brazil are not elected and have job security, [^2] By law, they acquire job stability after 2 years on the job.
-The goal is to insulate them from political pressure.
-promotion is the main form of career incentive.
-However, promotion from the Trial Court to the Appeals Court is rare and extremely unlikely in the short run, unlike career progression at earlier stages (Laneuville, 2024).
-We find that there is no gender conviction-rate gap for junior judges, while senior female judges convict domestic violence cases more frequently than male judges in the same career stage.
-This is an important avenue for future research because it could mean that the difference in perspective of early-career judges is not entirely reflected in the gender conviction-rate gap, potentially because of early-career judges’ desire to conform to the court’s established norms.
+Lastly, we discuss the consequences of the gender conviction-rate gap beyond the Trial Court.
+To do so, we analyze two types of consequences of a Trial Judge’s sentences.
+First, we look at the outcome of domestic violence cases in the Appeals Court, analyzing whether the Trial Judge’s gender impacts how frequently there is an appeal and how frequently their sentence is reversed.
+Second, we look at the future behavior of defendants in domestic violence cases, analyzing whether the Trial Judge’s gender impacts the defendants’ criminal recidivism after the final sentence.
+The impact of stricter female judges on appeals and ruling reversals in domestic violence cases is theoretically ambiguous.
+On the one hand, the positive gender conviction-rate gap may lead to female judges facing a larger reversal probability because sentences by stricter judges are more frequently reversed (Possebom, 2023).
+On the other hand, female judges may exert more effort (informational account) than male judges when analyzing domestic violence cases and, consequently, their rulings might have a smaller probability of being reversed.
+Importantly, in our empirical context, we find that these two forces balance each other and the Trial Judge’s gender has no impact on the outcome of domestic violence cases in the Appeals Court.
+For this reason, we conclude that the gender conviction-rate gap does not entail a higher cost in terms of Courts’ resources.
+The impact of stricter female judges on recidivism by defendants in domestic violence cases is also theoretically ambiguous.
+On the one hand, the positive gender conviction-rate gap may lead to female judges facing more recidivism because stricter judges tend to convict the type of defendants whose recidivism increases when punished (Acerenza, Possebom, and Sant’Anna, 2024).
+On the other hand, female judges may be able to influence defendants’ behaviors in other ways, such as directly signaling that violence against women is not tolerated (representational account).
+If the typical defendant in a domestic violence case is more likely to recidivate in a similar offense than to commit other types of crime, female judges may lead to lower recidivism.
+Importantly, in our empirical context, we find that these two forces balance each other and the Trial Judge’s gender has no impact on recidivism.
+For this reason, we conclude that the gender conviction-rate gap has no negative consequences in terms of future criminal behavior by defendants.
 This article contributes to the literature on judicial biases.
@@
 Within the studies focusing on in-group bias, many authors find evidence of judges making different decisions when they share a group identity with defendants or plaintiffs (Kruttschnitt and Savolainen, 2009; Boyd, Epstein, and Martin, 2010; Shayo and Zussman, 2011, 2017; Bielen and Grajzl, 2020; Cai et al., 2022; Chen et al., 2022).
-Similarly to our work, Kruttschnitt and Savolainen (2009), Boyd, Epstein, and Martin (2010) Bielen and Grajzl (2020), and Chen et al.
-(2022) analyze in-group gender bias in cases of uncommon crime types (rape and human trafficking) in Finland, federal appellate cases in the U.S., civil cases in Belgium, and for female offenders in Kenya, respectively.
-Our work differs from theirs in that we focus on a single common crime type (domestic violence) whose victims are mostly women.
+Similarly to our work, Kruttschnitt and Savolainen (2009), Boyd, Epstein, and Martin (2010), Bielen and Grajzl (2020), and Chen et al.
+(2022) analyze in-group gender bias in cases of uncommon crime types (rape and human trafficking) in Finland, federal appellate cases in the U.S., civil cases in Belgium, and in cases with female offenders in Kenya, respectively.
+Our work differs from theirs because we focus on a single common crime type (domestic violence) whose victims are mostly women.
 Furthermore, while Boyd, Epstein, and Martin (2010) finds that in-group bias in federal appellate cases in the U.S.
 is consistent with the informational account only, we find that in-group bias in criminal cases in Brazil is consistent with the representational and informational accounts.
+More importantly and differently from the previous studies, we not only document in-group bias by female judges in domestic violence cases but also analyze the consequences of this phenomenon beyond the Trial Court.
+We investigate the effects of the Trial Judge’s gender on the case’s outcomes at the Appeals Court and on the defendant’s future criminal behavior.
+In particular, we find that the gender conviction-rate gap has no negative consequences in terms of the use of Courts’ resources nor in terms of the defendants’ future criminal behavior.
 The paper is organized as follows.
 Section 2 describes our dataset, while Section 3 explains our empirical strategy.
-Moreover, Section 4 documents the gender conviction rate gap, while Section 5 discusses its driving forces.
-Finally, Section 6 argues that the difference in perspective of early-career judges may not be entirely reflected in the gender conviction-rate gap, and Section 7 concludes.
+Moreover, Section 4 documents the gender conviction-rate gap, while Section 5 discusses its driving forces.
+Section 6 analyzes the consequences of the gender conviction-rate gap beyond the Trial Court.
+Section 7 concludes.
 Data Description and Summary Statistics
@@
 We collect data from all criminal cases brought to the Justice Court System in the State of São Paulo, Brazil, between January 4th, 2010, and December 3rd, 2019.
-This dataset includes the date each case was assigned to a courtroom, the court district where the crime took place, the judge assigned to the case, the full text of the sentence, a list of the main events of the case, and the primary and secondary crime types associated to the case.
+This dataset includes the date each case was assigned to a courtroom, the court district where the crime took place, the judge assigned to the case, the full text of the sentence, a list of the main events of the case, the primary and secondary crime types associated to the case, information on whether the case was analyzed by the Appeals Court, and the full text of the Appeals Court’s ruling.
 The data also contain two pieces of information that are crucial for the construction of our sample, which we discuss in detail in Section 2.3.
@@
 We use this information to make sure that domestic violence cases are indeed randomly assigned within a district and not systematically assigned to (or reassigned to) judges who do not preside over cases of other types of crimes.
-We use the full-sentence text to construct our main outcomes.
-To do this, we adapt the classification algorithm developed by Possebom (2023) so that it classifies domestic violence case outcomes in a more precise manner.
-[^4] See Appendix A for an overview of our data-construction process.
-We classify sentences into one of six possible judicial decision categories.
-The first decision type is conviction.
-This variable equals one when the judge convicts at least one of the defendants listed in the case of at least one of the charges on the case.
-As a consequence of conviction, the defendant receives some type of punishment (e.g., incarceration, fine, or community service) and a criminal record.
-The second decision category is a non-prosecution agreement (NPA).
-This variable equals one when the defendant and the district attorney’s office strike a deal to exchange a criminal charge for a lighter punishment, such as a fine.
-Such agreements must be authorized by a judge and do not generate a criminal record.
-The third decision type is dismissal.
-A judge may dismiss a case if they consider it a less serious offense and the defendant fulfills a series of requirements.
-For example, the defendant may be forced to show up in court periodically for two years and be forbidden from going to bars.
-Similarly to NPAs, dismissed cases do not generate criminal records.
-The fourth decision category is acquittal. This variable equals one if the defendant is ruled “not guilty” on the charges listed.
-Lastly, we account for cases that do not come to a conclusion because either the defendant died before sentencing or because the criminal charges expired.
-In the latter situation, defendants can no longer be convicted according to the statute of limitations because too much time has elapsed since the day the crime was committed.
 We use the list of the case’s main events to measure two key case characteristics.
@@
 Sheriffs, who must have a Law degree in Brazil, may temporarily detain a defendant if they believe that the defendant poses a significant threat to society.
-The district attorney’s office defines each case’s primary and secondary crime prior to judge assignment. We construct three samples based on those crime types.
+We use the full-sentence text to construct our main outcome.
+To do this, we adapt the classification algorithm developed by Possebom (2023) so that it classifies domestic violence case outcomes in a more precise manner.
+[^3] See Appendix A for an overview of our data-construction process.
+Differently from Possebom (2023), our main outcome focuses exclusively on conviction decisions.
+This variable equals one when the judge convicts at least one of the defendants listed in the case of at least one of the charges on the case.
+As a consequence of conviction, the defendant receives some type of punishment (e.g., incarceration, fine, or community service) and a criminal record.
+We also collect information on three types of outcomes that may happen after the Trial Judge’s ruling.
+First, we measure whether the case was analyzed by the Appeals Court.
+This event happens if either the defense attorney or the district attorney disagrees with the Trial Judge’s ruling and appeals to a panel of three senior judges.
+Second, using the classification algorithm proposed by Possebom (2023), we analyze the full text of the Appeals Court’s ruling to measure whether the Trial Judge’s ruling was reversed or not.
+[^4] See Appendix A for an overview of our data-construction process.
+Third, we observe whether the defendant recidivated after their final sentence’s date and the type of the new criminal offense.
+To measure recidivism, we check whether the defendant’s name appears in any criminal case within two years after the final sentence date.
+[^5] See Appendix A for an overview of our data-construction process.
+Furthermore, the district attorney’s office defines each case’s primary and secondary crime type prior to judge assignment. We construct three samples based on this classification.
 First, our ‘domestic violence’ sample contains all cases with (i) domestic violence as the main subject or (ii) physical assault as the primary type and domestic violence as the secondary type.
@@
 11,340/2006) and increases the punishment severity of physical assault crimes associated with domestic violence.
-We define a ‘physical assault’ sample of cases whose primary type is physical assault.
-These cases are associated with other forms of assault such as street fights, bar fights, or football hooliganism.
+Second, our ‘physical assault’ sample contains cases whose primary type is physical assault.
+These cases are associated with other forms of assault, such as street fights, bar fights, or football hooliganism.
 We call ‘other physical assault’ the physical assault cases that do not involve domestic violence.
-Third, we define a ‘misdemeanors’ sample comprised of those crimes that carry a maximum incarceration sentence of less than four years.
-This threshold is relevant because, according to Brazilian Law, being convicted of these crime types leads to a criminal record, but not to incarceration.
-Sentences must be converted into fines or community service.
+Third, our ‘misdemeanors’ sample contains cases that carry a maximum incarceration sentence of less than four years.
+This threshold is relevant because, according to Brazilian Law, being convicted of these crime types leads to a criminal record but not to incarceration: sentences must be converted into fines or community service.
 As discussed by Possebom (2023) and Acerenza, Possebom, and Sant’Anna (2024), most misdemeanor crimes in Brazil are theft offenses.
@@
 We link this dataset to the criminal cases dataset by matching the judge’s name.
-The productivity data from the São Paulo Court of Justice website contains records on the total amount of work done by each judge in any given month and court district from January 2011 to December 2019.666For example, it measures the number of hearings or disposed cases per month for each judge in each court district.
+The productivity data from the São Paulo Court of Justice website contains records on the total amount of work done by each judge in any given month and court district from January 2011 to December 2019.777For example, it measures the number of hearings or disposed cases per month for each judge in each court district.
 We use this information to keep track of which judges are active in each court district-month pair.
@@
 Sample Definition and Summary Statistics
-Our final dataset is at the criminal-case level. It is crucial for our identification strategy that, for each case, the gender of the judge is randomly assigned and correctly identified.
+Our final dataset is at the criminal-case level.
+It is crucial for our identification strategy that the gender of the judge is randomly assigned and correctly identified for each case.
+To enforce these two characteristics, we adopt four steps.
 First, we restrict the dataset to cases that were randomly allocated to judges.
@@
 For this reason, we restrict the sample to district-quarter pairs without a courtroom specializing in domestic violence.
-Third, since we compare decisions made by female and male judges, we require the judge’s gender to be randomly allocated to cases.
+Third, since we compare decisions made by female and male judges, we require the judge’s gender to be randomly allocated to cases.101010 We measure judges’ gender using a two-step procedure.
+First, we apply the classification algorithm proposed by the R package genderBR, which uses the share of men and women with each name in the 2010 Brazilian Census.
+Since a few names are not commonly associated with a specific gender, we visit the São Paulo Court System’s website, check the photos of the remaining judges and manually classify their gender.
+The same classification procedure was used by Laneuville (2024).
 In other words, the ex-ante probability of a female judge being assigned to each case must be different from zero or one, which would not be the case in courts without judges of both genders.
@@
 To do this, we focus on cases allocated to judges between January 2011 and December 2019 and use our productivity dataset to remove all observations in which the assigned judge was not active at the court district and month of the assignment.
-[^9] Cases are randomly allocated to courtrooms and instead of directly to individual judges.
+[^11] Cases are randomly allocated to courtrooms instead of directly to individual judges.
 If a court district is understaffed at the time of case assignment and there is no presiding judge in the assigned courtroom, the system may record the case as being attributed to the next person who takes over the courtroom, which may take place at a later point in time.
-Using this final sample, we compute summary statistics for each possible case outcome and crime type.
-Table 1 displays case outcomes in the columns and crime types in the rows.
+Using this final sample, we compute summary statistics for cases’ outcomes and crime types.
+Table 1 displays cases’ outcomes in the columns and crime types in the rows.
 It also reports the share of female judges analyzing each type of crime.
-Domestic violence cases are more frequently acquitted than other misdemeanors or physical assault cases.
-As conviction rates are similar across types of crimes, this higher acquittal is due to fewer NPAs and dismissed cases.
+Domestic violence cases are more frequently convicted than other physical assault cases but are less frequently convicted than other misdemeanor cases.
 We also note that the share of female judges is similar across crime types and that there are 270 female judges and 425 male judges in our entire sample.
-Table 1: Case Outcomes and Share of Female Judges by Crime Type in %
-Crime Conviction NPA Acquittal Dismissal Expired Dead Female Obs
+Table 1: Cases’ Outcomes and Share of Female Judges by Crime Type in %
+Crime Conviction Female Obs
 Type Judge
-DV 32.5 1.1 51.4 4.3 8.8 2.0 32.1 4,358
-Other Mis 36.9 5.4 31.2 15.5 9.5 2.2 36.4 39,815
-Other PA 27.2 7.0 34.1 20.6 10.4 1.9 38.1 1,648
-  • Notes: The rows of this table indicate the different samples used in this paper: domestic violence cases (DV), misdemeanor cases unrelated to domestic violence (Mis), and physical assault cases that are not domestic violence (PA).
-The first six columns indicate the frequency of each type of case outcome.
+DV 32.5 32.1 4,358
+Other Mis 36.9 36.4 39,815
+Other PA 27.2 38.1 1,648
+  • Notes: The rows of this table indicate the different samples used in this paper: domestic violence cases (DV), misdemeanor cases unrelated to domestic violence (Other Mis), and physical assault cases that are not domestic violence (Other PA).
+The first column indicates the frequency of cases whose final ruling is conviction.
 The conviction dummy equals one when at least one defendant is convicted in at least one charge.
-The NPA dummy equals one when the defendant and the district attorney’s office sign an NPA.
-The dismissal dummy equals one if the case is dismissed.
-The acquittal dummy equals one if the defendant is found not guilty.
-The “expired” dummy equals one if at least one of the charges in the case has expired.
-The “Dead” dummy equals one if the defendant died before the sentence’s date.
-The rows do not sum to 100% because cases with multiple charges can have different outcomes for each charge.
-The second to last column shows the share of female judges in each sample while the last column shows the size of each sample.
+The second column shows the share of female judges in each sample, while the last column shows the size of each sample.
 Empirical Strategy and Validity Tests
 In this section, we explain the empirical strategy used to measure the gender conviction-rate gap and indirectly test the strategy’s validity.
-We measure the gender conviction rate gap for domestic violence cases by comparing the conviction probability for similar cases allocated to female and male judges.
+Empirical Strategy
+We measure the gender conviction-rate gap for domestic violence cases by comparing the conviction probability for similar cases allocated to female and male judges.
 We do this by running a linear regression model on the sample of domestic violence crimes:
 $\text{Convicted}_{it}=\beta\cdot\text{Female Judge}_{it}+\eta\cdot X_{it}+\text{Court District}_{i}\times\text{Quarter}_{t}+\varepsilon_{it}$ (1)
-where $\text{Convicted}_{it}$ indicates whether the cases resulted in a conviction, $\text{Female Judge}_{it}$ indicates whether the case was assigned to a female judge, $X_{it}$ are covariates including case characteristics (an indicator for whether the defendant was already in jail and an indicator for whether the police caught them committing the crime) and pre-determined judge characteristics (career stage, total experience, and experience in their current career stage as of the day the case was assigned), and $\text{Court District}_{i}\times\text{Quarter}_{t}$ is a full set of court district-quarter pair fixed effects.
+where $\text{Convicted}_{it}$ indicates whether the cases resulted in a conviction, $\text{Female Judge}_{it}$ indicates whether the case was assigned to a female judge, $X_{it}$ are covariates including pre-determined case characteristics (an indicator for whether the defendant was already in jail and an indicator for whether the police caught them committing the crime) and pre-determined judge characteristics (career stage, total experience, and experience in their current career stage as of the day the case was assigned), and “$\text{Court District}_{i}\times\text{Quarter}_{t}$” is a full set of court district-quarter pair fixed effects.
 Our parameter of interest, $\beta$, measures whether a female judge is more likely to convict than a male judge under the assumption that the judge’s gender is independent of the case’s potential outcome.
@@
 Even when our exogeneity assumption is valid, the $\beta$ parameter in Equation (1) may capture differences between male and female judges that are unrelated to their views on domestic violence.
-To measure whether this conviction-rate gap is due to different gender perspectives about this type of crime, we compare this domestic violence conviction gap against the gap for similar crime types.
+To measure whether this conviction-rate gap is due to different gender perspectives about this type of crime, we compare this domestic violence conviction-rate gap against this gap for similar crime types.
 Now, our sample consists of domestic violence cases and one type of comparable criminal case.
@@
 $\hskip 20.0pt+\eta\cdot X_{it}+\text{Court District}_{i}\times\text{Quarter}_{t}+\varepsilon_{it},$
-where $\text{Domestic Violence}_{it}$ indicates whether the case is associated with domestic violence.
+where “$\text{Domestic Violence}_{it}$” indicates whether the case is associated with domestic violence.
 We use two categories of comparable crime types: (i) offenses with similar sentences (misdemeanors) and (ii) offenses of a similar nature (physical assault).
@@
 This specification accounts for the extra strictness and social unacceptability of domestic violence offenses beyond that of crimes of comparable severity (coefficient $\beta_{2}$) and for the possibility that female judges are stricter than male judges in general (coefficient $\beta_{1}$).
-Consequently, this coefficient captures the part of the gender conviction-rate gap that is likely due to different gender perspectives about domestic violence cases specifically.
-For this reason, we interpret the $\beta_{3}$ coefficient as capturing a type of in-group bias according to the definition given by Shayo and Zussman (2011).
+Consequently, the $\beta_{3}$ coefficient captures the part of the gender conviction-rate gap that is likely due to different gender perspectives about domestic violence cases specifically.
+For this reason, we interpret this coefficient as capturing a type of in-group bias according to the definition given by Shayo and Zussman (2011).
 We do so because victims of domestic violence are overwhelmingly female, and members of this group (female judges) may act differently from members of the other group (male judges) when analyzing this type of crime in particular.
 This interpretation for the $\beta_{3}$ coefficient is valid if the judge’s gender is independent of the case’s potential outcome after conditioning on the type of crime.
-This exogeneity assumption is again valid according to Brazilian Law since cases are randomly allocated to judges within court districts and crime types are determined by the district attorney prior to judge assignment (Section 2.3).
+This exogeneity assumption is again valid according to Brazilian Law since cases are randomly allocated to judges within court districts, and crime types are determined by the district attorney prior to judge assignment (Section 2.3).
 We again test this assumption’s validity in Table 2.
+Validity Tests
 To test the validity of the identifying assumptions behind Equations (1) and (2), we measure the linear relationship between case characteristics and a dummy variable for an assigned female judge.
@@
 We find that, after including court district-quarter fixed effects, judge gender is uncorrelated with case characteristics such as whether the defendant was in jail before the case was assigned, whether the defendant was caught in the act by the police and whether the crime is a domestic violence offense.
-This result suggests that our identifying assumptions are valid.
+These results suggest that our identifying assumptions are valid.
 Table 2: Testing the Identification Assumption: Gender of the Judge and Case Characteristics
@@
 ${}^{***}p<0.01$; ${}^{**}p<0.05$; ${}^{*}p<0.1$
-  • Notes: Each cell represents the linear relationship between the gender of the judge randomly assigned to the case and a set of characteristics of the case.
+  • Notes: Each cell represents the linear relationship between the gender of the judge randomly assigned to the case and a set of case characteristics.
 Column (1) shows this relationship for the sample of all physical assault crimes, Column (2) shows this relationship for the sample of misdemeanors (crimes such that the maximum sentence does not exceed 4 years), and Column (3) shows this relationship for the sample of domestic violence crimes.
@@
 Measuring the Gender Conviction-Rate Gap
-We measure the gender conviction rate gap using two empirical strategies.
+We measure the gender conviction-rate gap using two empirical strategies.
 First, we estimate Equation (1) and find that female judges are more likely to convict in domestic violence cases than equally senior men.
-Second, we estimate Equation (2) and find that the gender conviction rate gap is small or nonexistent for misdemeanors and non-domestic violence physical assaults, whereas the gap becomes large and significant for domestic violence cases.
+Second, we estimate Equation (2) and find that the gender conviction-rate gap is small or nonexistent for misdemeanors and non-domestic violence physical assaults, whereas the gap becomes large and significant for domestic violence cases.
 Table 3 shows the effect of judge gender on the probability of conviction in domestic violence cases according to Equation (1).
@@
 All regressions include quarter-by-court district fixed effects.
-Since the results in Table 3 may capture differences between male and female judges beyond their gendered perspectives on domestic violence, we also estimate Equation (2) to compare the gender conviction rate gap for domestic violence cases against the gap for similar cases.
+Since the results in Table 3 may capture differences between male and female judges beyond their gendered perspectives on domestic violence, we also estimate Equation (2) to compare the gender conviction-rate gap for domestic violence cases against this gap for similar cases.
 Table 4 shows the estimated results associated with Equation (2).
@@
 The last three columns use the sample of physical assault cases, which are similar in nature to domestic violence cases as they entail physical harm to another person.
+Table 4 shows that the estimated coefficient on the female dummy ($\beta_{1}$) is small for the misdemeanors sample and close to zero for the physical assault sample.
+This suggests that the gender conviction-rate gap is small or nonexistent for comparable crimes that are not classified as involving domestic violence.
+More importantly, Table 4 shows that the coefficient on the interaction of judge gender and domestic violence ($\beta_{3}$) is statistically significant and positive in all regressions, regardless of included control variables.
+The effect is particularly strong for the physical assault sample (Columns (4)-(6)).
+This result suggests that the gender difference in conviction rates verified in Table 3 is neither fully explained by the severity or the nature of domestic violence offenses nor by female judges being stricter than male judges in all rulings.
+Consequently, the interaction coefficient captures the part of the gender conviction-rate gap that is specific to domestic violence cases and is likely due to different approaches to domestic violence cases between male and female judges.
+For this reason, this result suggests the existence of a type of in-group bias according to the definition given by Shayo and Zussman (2011).
+[^12] In Appendix B, we propose a simple economic model formalizing the definition of in-group bias.
+In the next section, we explore the driving forces behind this in-group bias, while, in Section 6, we investigate its consequences for the cases’ later stages in the Justice System and the defendants’ future behavior.
 Table 4: Effect of Judge Gender on the Probability of Conviction Comparing Domestic Violence Cases against Similar Crimes
@@
   • Notes: This table presents the estimated results of Equation (2).
-The first three columns compare differences in male and female judge conviction rates in domestic violence cases to the difference in conviction rates for misdemeanors.
-The last three columns compare differences in male and female judge conviction rates in domestic violence cases to the difference in conviction rates for other physical assault crimes.
+The first three columns compare differences in male and female judge conviction rates in domestic violence cases to the gender difference in conviction rates for misdemeanors.
+The last three columns compare differences in male and female judge conviction rates in domestic violence cases to the gender difference in conviction rates for other physical assault crimes.
 CR stands for Conviction Rate and DV stands for Domestic Violence.
-Table 4 shows that the estimated coefficient on the female dummy ($\beta_{1}$) is small for the misdemeanors sample and close to zero for the physical assault sample.
-This suggests that the gender conviction rate gap is small or nonexistent for comparable crimes that are not classified as involving domestic violence.
-More importantly, Table 4 shows that the coefficient on the interaction of judge gender and domestic violence ($\beta_{3}$) is statistically significant and positive in all regressions, regardless of included control variables.
-The effect is particularly strong for the physical assault sample (Columns (4)-(6)).
-This result suggests that the difference in conviction rates verified in Table 3 is neither fully explained by the severity or the nature of domestic violence offenses nor by female judges being tougher than male judges across the board.
-The interaction coefficient captures the part of the gender conviction rate gap that is specific to domestic violence cases and is likely due to different approaches to domestic violence cases between male and female judges.
-For this reason, this result suggests the existence of a type of in-group bias according to the definition given by Shayo and Zussman (2011).
-[^10] In Appendix C, we propose a simple economic model formalizing the definition of in-group bias.
-In the next section, we explore the driving forces behind this in-group bias.
-Understanding the Gender Conviction Rate Gap
+Understanding the Gender Conviction-Rate Gap
 In this section, we show two important findings about the conviction rate gender gap and its driving forces.
-The first is suggestive evidence that the gap is explained by judges’ identification with their own gender (Section 5.1).
+The first finding is suggestive evidence that the gap is explained by judges’ identification with their own gender (Section 5.1).
 We find that female judges are more likely than male judges to convict and simultaneously flag the case as an intimate partner violence offense, while female and male judges are equally likely to convict and not flag the case as an intimate partner violence offense.
-This result suggests that the representational account (Boyd, Epstein, and Martin, 2010) is one of the explanatory forces of the in-group bias documented in Table 4 because violence between relationship partners is highly impacted by patriarchal social norms that make gender identity more salient (Heise and Kotsadam, 2015; González and Rodríguez-Planas, 2020).
+This result suggests that the representational account (Boyd, Epstein, and Martin, 2010) is one of the explanatory forces of the in-group bias documented in Table 4 because violence between relationship partners is highly impacted by patriarchal social norms that make gender identity more salient (Heise and Kotsadam, 2015; Tur-Prats, 2019; González and Rodríguez-Planas, 2020).
 The second finding, discussed in Section 5.2, is that men and women respond differently to evidence presented in the same case.
-We present evidence that female judges are more likely than male judges to convict in cases where evidence is more tenuous and subject to interpretation, but not when the evidence is objectively incriminating.
+We show that female judges are more likely than male judges to convict in cases where evidence is more tenuous and subject to interpretation, but not when the evidence is objectively incriminating (i.e., when the defendant was caught red-handed).
 This result suggests that the informational account (Boyd, Epstein, and Martin, 2010) is one of the explanatory forces of the in-group bias documented in Table 4.
-In Appendix C, we propose a simple threshold-crossing model that formalizes the concepts of the representational and informational accounts.
-Female Identity as a Driver for Conviction Rates
+In Appendix B, we propose a simple threshold-crossing model that formalizes the concepts of the representational and informational accounts.
+Female Identity as a Driver for Larger Conviction Rates
 In this section, we show that the gender of the judge matters more in cases where patriarchal social norms are particularly relevant: intimate partner violence.
-We interpret this finding as indicative that a major reason for the gender conviction gap is that domestic violence cases make gender identity more salient to judges, suggesting that the representational account (Boyd, Epstein, and Martin, 2010) is one of the explanatory forces of the in-group bias documented in Section 4.
+We interpret this finding as indicative that a major reason for the gender conviction-rate gap is that domestic violence cases make gender identity more salient to judges, suggesting that the representational account (Boyd, Epstein, and Martin, 2010) is one of the explanatory forces of the in-group bias documented in Section 4 because female judges tend to protect member of their own group more intensely than male judges.
 We argue that domestic violence cases make gender identity more salient because sexist social norms and gender roles significantly affect the occurrence and punishment of intimate partner violence in multiple contexts.
 For instance, González and Rodríguez-Planas (2020) show that, among immigrant families, the incidence of intimate partner violence is positively affected by gender norms in the country of origin.
-Heise and Kotsadam (2015) show that the incidence of intimate partner violence in different countries is strongly associated with patriarchal social norms, such as male authority over women, justifications for wife beating, and laws and practices that disadvantage women in property ownership and other economic resources.
+Heise and Kotsadam (2015) show that the incidence of intimate partner violence in different countries is strongly associated with patriarchal social norms, such as male authority over women, justifications for wife beating, and laws and practices that disadvantage women in property ownership.
 Closer to our context, Perova and Reynolds (2017) show that police stations that specialize in domestic violence with specially trained officers significantly reduce the number of deaths of young women by intimate partner violence in metropolitan municipalities in Brazil.
@@
 The specific words we look for are “namorada,” “ex-namorada,” “esposa,” “ex-esposa,” “amásia,” “ex-amásia,” “companheira,” “ex-companheira,” “mulher,”, “ex-mulher,” “coabitação,” “união estável,” “casamento,” “era casado,” and “era casada.”
-We, then, test whether female judges are more likely to make a conviction decision in a domestic violence case and flag the case as being related to intimate partner violence after controlling for case characteristics.
-If there is a gender conviction gap in cases flagged as intimate partner violence but not in cases without such a flag, then a stronger identification with the victim may be driving the gender lenience gap found in Section 4.
-If the gender conviction rate gap is present regardless of explicit mentions of intimate partner violence, then there must be other forces driving this lenience gap.
-To implement this test, we interact the indicator for assault between intimate partners with the conviction indicator. This interaction term becomes our new outcome variable in Equation (1).
+We, then, test whether female judges are more likely to make a conviction decision in a domestic violence case and simultaneously flag the case as being related to intimate partner violence after controlling for case characteristics.
+If there is a gender conviction-rate gap in cases flagged as intimate partner violence but not in cases without such a flag, then a stronger identification with the victim may be driving the gender lenience gap found in Section 4.
+If the gender conviction-rate gap is present regardless of explicit mentions to intimate partner violence, then there must be other forces driving this lenience gap.
+To implement this test, we interact the “intimate partner violence” indicator with the conviction indicator. This interaction term becomes our new outcome variable in Equation (1).
 Table 5 shows the results of these regressions.
-In Column (1), the outcome is a dummy variable indicating whether the case included intimate partner violence and resulted in a conviction.
-In Column (2), the outcome is a dummy variable indicating whether the case did not include intimate partner violence and also resulted in a conviction.
+In Column (1), the outcome is a dummy variable indicating whether the judge’s ruling mentioned intimate partner violence and reached a conviction decision.
+In Column (2), the outcome is a dummy variable indicating whether the judge’s ruling did not mention intimate partner violence but reached a conviction decision.
 In Column (3), the outcome is a dummy variable indicating whether the case resulted in a conviction, replicating the results from Table 3 for convenience.
-We find that the probability of an intimate partner violence case resulting in conviction is 37.5% (9 p.p.) higher when the judge is female (Column (1)).
-However, the probability of a case not involving intimate partner violence resulting in conviction is the same regardless of the gender of the judge (Column (2)).
-These results suggest that the gender conviction rate gap is higher in cases where a female judge may identify more strongly with the victim compared to a male judge.
+We find that female judges are 37.5% (9 p.p.) more likely than male judges to convict the defendant and simultaneously flag the case as intimate partner violence (Column (1)).
+However, female judges are as likely as male judges to convict the defendant while not flagging the case as intimate partner violence (Column (2)).
+These results suggest that the gender conviction-rate gap is higher in cases where a female judge may identify more strongly with the victim compared to a male judge.
 Table 5: Effect of Judge Gender on the Probability of Conviction and the Salience of Intimate Partner Violence
@@
 ${}^{***}p<0.01$; ${}^{**}p<0.05$; ${}^{*}p<0.1$
-  • The regressions in this table refer to outcomes for the sample of domestic violence crimes.
+  • The regressions in this table refer to outcomes for the sample of domestic violence crimes only.
 The outcome in Column (1) is whether the case involved intimate partner violence and led to a conviction.
@@
 (2022) show that judges analyzing divorce cases in China favor plaintiffs of their own gender more in areas where social norms are more conservative.
-In our case, we show that the gender gap is much more intense in cases where the form of domestic violence is more stereotypical.
+In our case, we show that the gender conviction-rate gap is much more intense in cases where the form of domestic violence is more stereotypical.
 [^14] Note that most domestic violence cases against women are committed by partners and ex-partners (United Nations Office on Drugs and Crime, 2023).
-Consequently, the representational account (Boyd, Epstein, and Martin, 2010) is a plausible driving forcing of the in-group bias documented in Section 4.
+Consequently, the representational account (Boyd, Epstein, and Martin, 2010) is a plausible driving forcing of the in-group bias documented in Section 4 because female judges tend to protect members of their own group more intensely than male judges.
 Male and Female Judges Analyze Evidence Differently
 In this section, we show that female judges are more likely than male judges to convict in domestic violence cases where evidence is more tenuous and subject to interpretation, but not when the evidence is objectively incriminating.
-We interpret this result as suggesting that the informational account (Boyd, Epstein, and Martin, 2010) is one of the explanatory forces of the in-group bias documented in Table 4.
+We interpret this result as suggesting that the informational account (Boyd, Epstein, and Martin, 2010) is one of the explanatory forces of the in-group bias documented in Table 4 because female judges process the information contained in domestic violence cases differently from their male peers.
 Domestic violence cases require complex evidentiary rulings.
@@
 Since such testimony is inherently subjective, female and male judges may interpret testimony evidence differently.
-Our dataset allows us to investigate whether the gender conviction rate gap is stronger in cases where the evidence set is less clearly incriminating.
+Our dataset allows us to investigate whether the gender conviction-rate gap is stronger in cases where the evidence set is less clearly incriminating.
 For each case, we observe whether the defendant was arrested because they were caught in the act of committing the crime (i.e., in flagrante delicto).
-These cases have more objective evidence, often including testimony from police officers who witnessed the offense.
-If the gender conviction rate gap is stronger in cases where the defendant was not caught red-handed, then this may be because female and male judges interpret more subjective pieces of evidence differently (informational account).
-If the gender conviction rate gap is similar across domestic violence cases regardless of the type of evidence presented, then there must be other forces driving this lenience gap.
-To measure the gender conviction rate gap for cases with different evidence types, we re-estimate Equations (1) and (2) using two sub-samples.
-The first restricts our sample to cases where the defendant was caught red-handed, while the second is restricted to cases where the defendant was not caught in flagrante delicto.
+These cases have more objective evidence, often including testimony from police officers who directly witnessed the offense.
+If the gender conviction-rate gap is stronger in cases where the defendant was not caught red-handed, then the gender leniency gap may be because female and male judges interpret more subjective pieces of evidence differently (informational account).
+If the gender conviction-rate gap is similar across domestic violence cases regardless of the type of evidence presented, then there must be other forces driving this lenience gap.
+To measure the gender conviction-rate gap for cases with different evidence types, we re-estimate Equations (1) and (2) using two sub-samples.
+The first regression restricts our sample to cases where the defendant was caught red-handed, while the second regression restricts our sample to cases where the defendant was not caught in flagrante delicto.
 Table 6 reports regression results for the two different samples.
-The odd columns include cases where the defendant was caught red-handed, while the even columns include cases where this objective evidence was not presented.
+The odd columns include cases where the defendant was caught red-handed, while the even columns include cases where this objective evidence was not present.
 Columns (1) and (2) report the results associated with Equation (1), while Columns (3)-(6) report the results associated with the interaction model described in Equation (2).
@@
   • Notes: This table compares the determinants of the probability of conviction.
-Cases are split into a sub-sample in which the defendant was caught red-handed and a sample in which the defendant was not, requiring the justice system to rely on less objective evidence.
-The first two columns are constructed using the sample of domestic violence cases.
-Columns 3 and 4 use the sample of misdemeanors.
-Columns 5 and 6 use the sample of physical assault crimes.
+Cases are split into a sub-sample in which the defendant was caught red-handed (odd columns) and a sample in which the defendant was not caught red-handed (even columns), requiring the justice system to rely on less objective evidence.
+The first two columns are constructed using the sample of domestic violence cases only.
+Columns (3) and (4) use the sample of misdemeanors.
+Columns (5) and (6) use the sample of physical assault crimes.
 CR stands for Conviction Rate and DV stands for Domestic Violence.
-We find that a gender conviction rate gap exists only in those cases where the defendant was not caught red-handed.
+All regressions include quarter-district fixed effects and a full set of controls for judge career stages.
+We find that a gender conviction-rate gap exists only in those cases where the defendant was not caught red-handed.
 We only find significant results for the female coefficient ($\beta$) in Column (2) and the interaction coefficient ($\beta_{3}$) in Column (6).
@@
 These different interpretations of more subjective cases are consistent with the informational account being one possible driving force of the in-group bias documented in Section 4.
-Discussion: Career Incentives and a Potential Underestimation of the Gap
-In this section, we analyze the association between the strength of career incentives and the size of the conviction gender gap.
-We show that male and female judges differ more intensely at the end of their careers.
-This corresponds with the period where they face the lowest probability of upcoming performance evaluation for promotion to the Appeals Court.
-[^13] See Appendix B for more information on the proportion of cases analyzed by judges who apply for promotion in the near future Career pressures may lead early-career male and female judges to act more similarly, possibly in response to feedback from Appeals Court magistrates, who are overwhelmingly male (Laneuville, 2024).
-Since Brazilian judges cannot be fired and must receive the same salary, promotion is the main career incentive faced by magistrates in our sample.
-Judges are unelected officials who may occupy four positions during their careers in the Trial Court (substitute, initial, intermediary, and final).
-At the final stage of their careers, they may be promoted to the Appeals Court.
-To be promoted to the next career stage, Trial Court Judges must apply for promotion opportunities and be approved by a vote of the Appeals Court Judges.
-Since very few judges are promoted to the Appeals Court (Laneuville, 2024), career incentives are stronger for trial judges who are not at the final stage of the Trial Court.
-[^14] Appendix B provides descriptive statistics showing that final-stage judges face weaker career incentives.
-Since adopting different sentencing criteria from peer judges may hamper chances for promotion, we investigate whether judges who face weak career incentives are more likely to differ from their colleagues, and whether these differences correspond with judge gender.
-If we observe a larger gender conviction gap for final-stage judges, then weaker career incentives may explain the gender leniency gap in domestic violence cases documented in Section 4.
-If judges in all career stages have similar gender conviction rates, then there must be other forces driving this lenience gap.
-To measure the gender conviction gap for judges in different career stages, we re-estimate Equations (1) and (2) using two sub-samples.
-The first restricts our sample to cases assigned to final-stage judges, while the second restricts the sample to cases assigned to judges in earlier stages of the Trial Court.
-Table 7 presents the estimated coefficients of Equations (1) and (2) for these two sub-samples.
-Results for the sample of final-stage judges are provided in the odd columns.
-Results for the sample of early-stage judges are provided in the even columns.
-The header indicates whether the sample contains only domestic violence cases, misdemeanor cases, or physical assault cases.
-Table 7: Effect of Judge Gender on the Probability of Conviction by Judge Career Stage
-Sample: Sample: Sample:
-Domestic Violence Misdemeanors Physical Assault
-Final Early Final Early Final Early
+Consequences of the Gender Conviction-Rate Gap
+A gender conviction-rate gap for domestic violence cases, as found in Section 4, may have consequences for the cases’ later stages in the Justice System and the defendants’ future behavior.
+In particular, defense attorneys, district attorneys, Appeals Court judges, and defendants may change their behavior when facing a female judge in a domestic violence case.
+In Section 6.1, we investigate whether stricter female judges may lead to more appeals and ruling reversals in domestic violence cases.
+Although we may expect positive gender appeal-rate and reversal-rate gaps due to the fact that stricter judges are more frequently reversed by the Appeals Court (Possebom, 2023, Figure 1), we find that female judges’ rulings in domestic violence cases have similar outcomes in the Appeals Court when compared to male judges’ rulings.
+In Section 6.2, we investigate whether stricter female judges may lead to more recidivism by defendants in domestic violence cases.
+Since stricter female judges might be more likely to convict the type of defendants whose recidivism increases when punished (Acerenza, Possebom, and Sant’Anna, 2024), we may expect a positive gender recidivism gap.
+Despite this expectation, we find that female and male judges face similar recidivism probabilities in their domestic violence cases.
+Consequences of the Gender Conviction-Rate Gap at the Appeals Court
+A gender conviction-rate gap for domestic violence cases may impact how the defense attorneys, district attorneys and Appeals Court judges treat domestic violence cases after the Trial Judge has issued a ruling.
+In this section, we investigate whether these agents change their behavior when facing a female judge in a domestic violence case.
+On the one hand, the phenomenon of stricter female judges for domestic violence cases may lead to more appeals and reversals of trial judges’ rulings.
+This positive gender gap may be expected for two reasons.
+First, Possebom (2023, Figure 1 and Table J.1) finds that rulings by strict judges are more frequently reversed by the Appeals Court than rulings by more lenient judges, partially because defense attorneys are more likely to appeal than district attorneys.
+Combining this result with our finding that female judges are stricter than male judges when analyzing domestic violence cases, we may expect a positive gender gap in the probability of domestic violence cases being analyzed by the Appeals Court and in the probability of domestic violence rulings being reversed by the Appeals Court.
+Second, Laneuville (2024, Table 3) states that 93.55% of Appeals Court judges who were active between 2007 and 2022 were men.
+If we assume that judges of both genders behave similarly at the Trial Court and the Appeals Court, our results suggest that male Appeals Court judges may be more lenient when analyzing domestic violence cases and reverse Trial Judges’ conviction rulings more frequently.
+Consequently, we may expect a positive gender gap in the probability of domestic violence rulings being reversed by the Appeals Court.
+On the other hand, the informational account for the gender conviction-rate gap in domestic violence cases (Section 5.2) may lead to a smaller probability of female judges having their rulings reversed by the Appeals Court.
+According to the informational account of the in-group bias, female judges analyze domestic violence evidence in a different way, possibly implying that they write more complete or careful rulings.
+When we combine this possibility with the fact that the Appeals Court focuses more on the ruling’s form than the case’s substance, we may expect a negative gender gap in the probability of domestic violence rulings being reversed by the Appeals Court.
+We investigate which of these two forces dominates by analyzing if there is a gender gap in the probability of domestic violence cases being analyzed by the Appeals Court and in the probability of domestic violence rulings being reversed by the Appeals Court.
+To do so, we re-estimate Equations (1) and (2) using two new outcome variables.
+First, our left-hand side variable is an indicator that equals 1 if there is an appeal process associated with the case and 0 otherwise.
+Second, our outcome variable is an indicator that equals 1 if the Appeals Court reversed the Trial Judge’s ruling and 0 otherwise.
+To avoid sample selection issues, note that the “zero” group includes all cases that were not analyzed by the Appeals Court.
+Table 7 shows the estimated coefficients of Equations (1) and (2) when we use our appeal and reversal indicators as outcome variables.
+The left-hand side variable is the appeal indicator in the odd columns, while it is the reversal indicator in the even columns.
+The first two columns are constructed using the sample of domestic violence cases only (Equation (1)).
+Columns (3) and (4) use the sample of misdemeanors (Equation (2)), while Columns (5) and (6) use the sample of physical assault crimes (Equation (2)).
+All regressions include quarter-district fixed effects and a full set of controls for case characteristics and judge career stages.
+There are three results worth highlighting.
+First, the sign and significance of the female coefficient in Columns (3) and (5) show that female judges are slightly more likely to be appealed in misdemeanor and physical assault cases, respectively.
+This finding is not surprising because female judges are stricter than male judges (Column (3) in Table 4), and defense attorneys appeal more frequently than district attorneys (Possebom, 2023, Table J.1).
+Second, because female coefficients are statistically null in Columns (2), (4) and (6), we find that, even though female trial judges are more frequently appealed, there is no gender difference in the reversal probability.
+Third and more importantly, the coefficients associated with the consequences of the in-group bias (“female” in Columns (1) and (2) and “DV x Female” in Columns (3)-(6)) are not statistically significant.
+[^15] Remember that the in-group bias captures the part of the gender conviction-rate gap that is specific to domestic violence cases.
+The null result in the “female” coefficient in Columns (1) and (2) means that female and male trial judges are equally likely to be appealed and reversed in domestic violence cases.
+The “DV x Female” coefficient in Columns (3)-(6) means that the gender gap in the probability of being appealed and reversed is not significantly more intense in domestic violence cases than in other comparable cases.
+These null results suggest that the forces driving the gender appeal-rate and reversal-rate gaps upwards and downwards compensate each other.
+In other words, even though female judges are stricter than male judges when analyzing domestic violence cases, their rulings are complete and careful enough so that they do not generate more reversal decisions by the Appeals Court.
+Table 7: Effect of Judge Gender on the Probability of a Case Being Analyzed by the Appeals Court and of a Trial Sentence Being Reversed by the Appeals Court
+Sample: DV only Sample: Misdemeanors Sample: Physical Assault
+Appealed Reversed Appealed Reversed Appealed Reversed
 (1) (2) (3) (4) (5) (6)
-Fem $0.23^{***}$ $-0.03$ $0.06^{***}$ $-0.00$ $0.09^{**}$ $-0.13^{*}$
+Female $0.04$ $0.01$ $0.02^{**}$ $0.00$ $0.05^{*}$ $0.01$
 $(0.03)$ $(0.01)$ $(0.005)$ $(0.004)$ $(0.03)$ $(0.01)$
-DV $-0.08^{***}$ $0.02$ $0.04$ $0.06$
+DV $-0.09^{***}$ $-0.03^{***}$ $0.04^{*}$ $0.01$
 $(0.01)$ $(0.01)$ $(0.02)$ $(0.01)$
-DV x Fem $0.09^{***}$ $-0.02$ $0.10^{*}$ $0.11$
+DV x Female $0.01$ $0.01$ $-0.01$ $0.00$
 $(0.02)$ $(0.01)$ $(0.04)$ $(0.01)$
-Men’s CR (Non-DV) - - 0.36 0.41 0.24 0.37
-Men’s CR (All) $0.26$ $0.43$ $0.35$ $0.50$ $0.25$ $0.50$
+Men’s Average (Non-DV) $0.37$ $0.05$ $0.17$ $0.03$
+Men’s Average (All) $0.29$ $0.03$ $0.37$ $0.05$ $0.22$ $0.03$
 Quarter-District FE $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+Career Controls $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 Case Controls $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-Num. obs. $2,609$ $1,749$ $25,639$ $18,534$ $3,499$ $2,507$
+Num. obs. $3,668$ $3,668$ $46,251$ $46,251$ $5,868$ $5,868$
 ${}^{***}p<0.01$; ${}^{**}p<0.05$; ${}^{*}p<0.1$
-  • Notes: This table shows results of regressing the conviction indicator on the indicator variables for female judges, domestic violence cases, and the interaction female x domestic violence for six different samples.
-Column (1) shows the results for the domestic violence cases assigned to final-stage judges.
-Column (2) shows the results for domestic violence cases assigned to early-stage judges.
-Column (3) shows the results for misdemeanor cases assigned to final-stage judges.
-Column (4) shows the results for misdemeanor cases assigned to early-stage judges.
-Column (5) shows the results for physical assault cases assigned to final-stage judges.
-Column (6) shows the results for cases assigned to early-stage judges.
-All regressions include quarter x court district fixed effects.
-CR stands for Conviction Rate and DV stands for Domestic Violence.
-We find that the gender conviction rate gap is much stronger for final-stage judges than for early-stage judges.
-In the sample of domestic violence cases, final-stage female judges are almost twice as likely to convict as their male peers (23 p.p.
-larger), while early-stage judges have similar conviction rates across genders.
-We reach qualitatively similar conclusions when analyzing the interaction coefficients in the misdemeanors and other physical assault samples.
-Hence, the bulk of the gender gap discussed in this paper comes from judges at the end of their careers who face lower pressure to perform.
-If career pressures affect judges’ verdicts, then early-career judges of different genders may have different views on domestic violence than their conviction rates might suggest.
+  • Notes: This table shows the results of Equations (1) and (2) using two new outcome variables.
+In the odd columns, the outcome variable is binary and indicates whether the case was analyzed by the Appeals Court, while, in the even columns, the outcome variable is binary and indicates whether the trial judge’s sentence was reversed by the Appeals Court.
+The first two columns are constructed using the sample of domestic violence cases (Equation (1)).
+Columns (3) and (4) use the sample of misdemeanors (Equation (2)).
+Columns (5) and (6) use the sample of physical assault crimes (Equation (2)).
+DV stands for Domestic Violence.
+All regressions include quarter-district fixed effects and a full set of controls for case characteristics and judge career stages.
+Consequences of the Gender Conviction-Rate Gap for the Defendants’ Recidivism Behavior
+A gender conviction-rate gap for domestic violence cases may impact how defendants behave after their sentences.
+In this section, we investigate whether these agents change their behavior when facing a female judge in a domestic violence case.
+On the one hand, the phenomenon of stricter female judges for domestic violence cases may increase the probability of recidivism. This positive gender gap may be expected for two reasons.
+First, Acerenza, Possebom, and Sant’Anna (2024) analyzes misdemeanor cases in São Paulo during our sample period and finds that punishment increases two-year recidivism for agents who would be punished only by strict judges (type A defendant) while it decreases two-year recidivism for agents who would be punished by most judges, including the lenient judges (type B defendant).
+In our context, we find that female judges are stricter than male judges (Table 4, Column (3)), especially for domestic violence cases (Table 4, Columns (3) and (6)).
+Moreover, if we assume that female judges are stricter because they have a harsher bar for conviction (instead of having more information about domestic violence cases), then we expect female judges to convict more agents who are more likely to recidivate when punished (type A defendant) while being similarly likely to convict type B defendants.
+Consequently, we may expect a positive gender gap in the probability of a defendant recidivating.
+Second, domestic violence victims may be encouraged by conviction decisions to accuse their aggressors if they recidivate by committing a new domestic violence offense.
+If this is the case, we may expect a positive gender gap in the recidivism probability because female judges are stricter than male judges when analyzing domestic violence cases (Section 4).
+To avoid this mechanical increase in the recidivism probability, we also analyze the effect of a female judge on the probability of committing a new crime that is not a domestic violence offense.
+On the other hand, the representational account for the gender conviction-rate gap in domestic violence cases (Section 5.1) may lead to a smaller probability of recidivism.
+For example, female judges may be able to directly influence the behavior of the defendants by better signaling that violence against women is not tolerated.
+If the typical defendant in a domestic violence case is more likely to recidivate in a similar offense than to commit other types of crime, female judges may lead to lower recidivism.
+Consequently, we may expect a negative gender gap in the recidivism probability.
+We investigate which of these two forces dominates by analyzing if there is a gender gap in the probability of a defendant recidivating within two years of the case’s final sentence.
+To do so, we re-estimate Equations (1) and (2) using two new outcome variables.
+First, our left-hand side variable is an indicator that equals 1 if the defendant commits any new criminal offense within two years of their final sentence and 0 otherwise.
+Second, our outcome variable is an indicator that equals 1 if the defendant commits any new non-domestic violence offense within two years of their final sentence and 0 otherwise.
+Table 8 shows the estimated coefficients of Equations (1) and (2) when we use these two measures of recidivism as outcome variables.
+The left-hand side variable is the any-recidivism indicator in the odd columns, while it is the non-domestic violence recidivism indicator in the even columns.
+The first two columns are constructed using the sample of domestic violence cases only (Equation (1)).
+Columns (3) and (4) use the sample of misdemeanors (Equation (2)), while Columns (5) and (6) use the sample of physical assault crimes (Equation (2)).
+All regressions include quarter-district fixed effects and a full set of controls for case characteristics and judge career stages.
+Note that the coefficients associated with the consequences of the in-group bias (“female” in Columns (1) and (2) and “DV x Female” in Columns (3)-(6)) are not statistically significant.
+[^16] Remember that the in-group bias captures the part of the gender conviction-rate gap that is specific to domestic violence cases.
+We also highlight that the relevant point estimates in Columns (1)-(4) are either small or negative.
+These results suggest that the forces driving the gender recidivism gap upwards and downwards compensate each other.
+In other words, even though female judges are stricter than male judges when analyzing domestic violence cases, their stricter rulings are not criminogenic.
+Table 8: Effect of Judge Gender on the Probability of a Defendant Recidivating
+Sample: DV only Sample: Misdemeanors Sample: Physical Assault
+Any Excluding DV Any Excluding DV Any Excluding DV
+(1) (2) (3) (4) (5) (6)
+Female $0.01$ $-0.02$ $0.01$ $0.00$ $-0.02$ $-0.04$
+$(0.04)$ $(0.04)$ $(0.01)$ $(0.01)$ $(0.05)$ $(0.05)$
+DV $-0.02$ $-0.06^{***}$ $0.02$ $-0.01$
+$(0.02)$ $(0.01)$ $(0.04)$ $(0.04)$
+DV x Female $0.01$ $0.00$ $0.04$ $0.03$
+$(0.02)$ $(0.02)$ $(0.06)$ $(0.05)$
+Men’s Average (Non-DV) $0.35$ $0.35$ $0.25$ $0.24$
+Men’s Average (All) $0.32$ $0.28$ $0.35$ $0.34$ $0.30$ $0.27$
+Quarter-District FE $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+Career Controls $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+Case Controls $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+Num. obs. $2,665$ $2,665$ $21,319$ $21,319$ $3,467$ $3,467$
+${}^{***}p<0.01$; ${}^{**}p<0.05$; ${}^{*}p<0.1$
+  • Notes: This table shows the results of Equations (1) and (2) using two new outcome variables.
+In the odd columns, the outcome variable is binary and indicates whether the case’s defendant was prosecuted for any crime after their final sentence’s date, while, in the even columns, the outcome variable is binary and indicates whether the case’s defendant was prosecuted for any crime that is not a domestic violence offense.
+The first two columns are constructed using the sample of domestic violence cases (Equation (1)).
+Columns (3) and (4) use the sample of misdemeanors (Equation (2)).
+Columns (5) and (6) use the sample of physical assault crimes (Equation (2)).
+DV stands for Domestic Violence.
+All regressions include quarter-district fixed effects and a full set of controls for case characteristics and judge career stages.
 Conclusion
-In this paper, we provide evidence that cases assigned to female judges are significantly more likely to result in conviction than those assigned to male judges with similar career paths.
-This pattern is, to a large extent, due to the reasons that are specific to domestic violence.
+In this paper, we provide evidence that domestic violence cases assigned to female judges are significantly more likely to result in conviction than those assigned to male judges with similar career paths.
+This pattern is, to a large extent, due to the reasons that are specific to domestic violence cases.
 The gender difference in conviction rates is stronger for cases involving domestic violence than for crimes that are similar in nature (physical assault) and sentence length (misdemeanors), suggesting the existence of a type of in-group bias as defined by Shayo and Zussman (2011).
-Furthermore, we find suggestive evidence for two major drivers of the conviction gender gap in domestic violence.
-The first is that female judges are stricter than men in cases of intimate partner violence but equally strict for cases of other domestic violence offenses.
-We claim this is evidence of identification with one’s own gender, as perception of intimate partner violence is highly affected by patriarchal social norms, as opposed to other forms of violence where this might not be the case.
+Furthermore, we find suggestive evidence for two major drivers of the gender conviction-rate gap in domestic violence.
+The first driver is that female judges are stricter than men in cases of intimate partner violence but equally strict in cases of other types of domestic violence.
+We claim this finding shows the relevance of identification with one’s own gender, as the perception of intimate partner violence is highly affected by patriarchal social norms, while other forms of violence are not affected by gender stereotypes.
 This result is consistent with the representational account as defined by Boyd, Epstein, and Martin (2010).
 The second driver comes from gender differences in the way judges analyze evidence.
-There is no gender gap when the defendant is caught red-handed by the police, making the evidence presented clear and objective, but the gap is very strong when there is more room for subjectivity in the analysis.
+There is no gender gap when the defendant is caught red-handed by the police (clear and objective piece of evidence), but the gap is very strong when there is more room for subjectivity in the analysis.
 This result is consistent with the informational account as defined by Boyd, Epstein, and Martin (2010).
-Finally, we identify an association between a larger gender gap and weaker career pressures for more senior judges in the final stage of the trial court.
-We believe this to be an important avenue for future research.
+Lastly, we find that the gender conviction-rate gap has no impact on the probability of appeals, ruling reversals or recidivism for female judges in domestic violence cases.
+Consequently, we conclude that there is no evidence that the gender conviction-rate gap affects the use of Courts’ resources or the defendants’ future criminal behavior.
 Supporting Information
@@
 For additional legal definitions, check Appendix I.2 by Possebom (2023).
-The final dataset was created from two initial datasets.
+The final dataset was created from three initial datasets.
   1. 1.
@@
     CJPG (“Consulta de Julgados de Primeiro Grau”): It contains information about the last decision made by a trial judge in all criminal cases in TJ-SP between 2010 and 2019.
+  3. 3.
+    CPOSG (“Consulta de Processos de Segundo Grau”): It contains information about all appealing criminal cases in TJ-SP between 2010 and 2019.
 Starting from the CPOPG dataset, we implement the following steps.
@@
 We then validate this algorithm using our three validation samples and find that it correctly classifies 97.9% of the overall cases, 97.9% of the domestic violence cases, and 95.3% of the physical assault cases.
-Given this high success rate, we use the L1-Regularized Logistic Regression algorithm to define the treatment variable in our full sample.
-Having designed the above algorithm, we use it to define the trial judge’s treatment variable $Y$ in the full sample.
+Given this high success rate, we use the L1-Regularized Logistic Regression algorithm to classify which trial judge rulings concluded that the “defendant is guilty” in our full sample.
+Having designed the above algorithm, we use it to define whether the trial judge reached a conviction ruling, which is our main outcome variable.
 First, we find which defendants died during their trials and drop them from our sample.
@@
 Moreover, we use the trained L1-regularized Logistic Regression algorithm to classify the remaining cases into the categories “defendant is guilty” and “defendant was acquitted”.
-Appendix B Final-stage Judges Face Weaker Career Incentives
-In Section 6, we find that only final-stage judges present a gender conviction rate gap for domestic violence cases and conclude that weaker career incentives for senior judges are one possible force driving the gender leniency gap.
-This interpretation relies on the assumption that early-stage judges face stronger career incentives than final-stage judges.
-The current appendix verifies the plausibility of this assumption.
-To do so, we estimate the percentage of cases in which the assigned judge applied for a promotion opportunity within two years after the case assignment.
-Table B.1 shows these probabilities by crime type and judge characteristics.
-The first row focuses on the set of cases in which the assigned judge was not in the final stage of the Trial Court System, while the second row focuses on the set of cases in which the assigned judge was in the final stage of the Trial Court System.
-Odd columns restrict the sample to cases attributed to female judges while even columns restrict the sample to cases attributed to male judges.
-The header indicates the sample used to estimate these probabilities.
-Table B.1: Percentage of Cases in which the Assigned Judge Applied for a Promotion Opportunity within Two Years by Crime Type and Judge Characteristics
-Sample: Sample: Sample:
-Judge Domestic Violence Physical Assault Misdemeanors
-Career Stage Female Male Female Male Female Male
-(1) (2) (3) (4) (5) (6)
-Early 46.7% 50.9% 46.2% 49.2% 41.4% 48.4%
-Final 0.0% 3.9% 0.0% 3.7% 0.0% 2.0%
-  • Notes: This table shows the probability of each type of case being assigned to a judge who applied for a promotion opportunity within two years of the day the case was allocated to them.
-The first row shows this statistic for the set of cases that were attributed to judges who were not in the final stage of the Trial Court System (i.e., substitute judges or magistrates in the initial or intermediate stages).
-The second row shows this statistic for the set of crimes attributed to judges in the final stage of the Trial Court System.
-Odd columns restrict the sample to cases attributed to female judges.
-Even columns restrict the sample to cases attributed to male judges.
-Columns (1)-(2) refer to cases in the domestic violence sample.
-Columns (3)-(4) use the sample of physical assault cases.
-Columns (5)-(6) refer to cases from the misdemeanors sample.
-We find that final-stage judges are much less likely to apply for promotion opportunities than early-stage judges.
-This result suggests that early-stage judges are more eager to be promoted and may face stronger career incentives than final-stage judges.
-Consequently, the assumption required to interpret weaker career incentives as one of the driving forces of the gender conviction rate gap is plausible.
-Lastly, we briefly discuss whether the different conviction criteria adopted by male and female judges may have career consequences for the trial judges.
-As explained by Laneuville (2024), Appeals Court Judges vote on which trial judge to promote.
-Consequently, adopting sentencing criteria that lead to more appeal processes and reversed decisions may impact the judge’s career prospects.
-To analyze the association between appeal processes, decision reversals, judges’ genders, and judges’ decisions, we regress an outcome variable indicating whether the trial judge’s decision was appealed or reversed on an indicator of a female judge, an indicator of a conviction decision, the interaction of the last two indicators, and district-by-quarter fixed effects.
-Table B.2 shows the results for the regression whose outcome variable indicates whether the trial judge’s decision was appealed while Table B.3 shows the results for the regression whose outcome variable indicates whether the trial judge’s decision was reversed.
-Table B.2: Association between Conviction Sentences and Appealing of Judges’ Decisions on Domestic Violence Cases
-Final Early
-Appealed Appealed Appealed Appealed
-(1) (2) (3) (4)
-Fem $0.17^{***}$ $0.09^{***}$ $-0.01$ $0.03$
-$(0.03)$ $(0.03)$ $(0.06)$ $(0.05)$
-Convicted $0.45^{***}$ $0.41^{***}$
-$(0.03)$ $(0.04)$
-Fem x Convicted $-0.06$ $-0.09$
-$(0.05)$ $(0.07)$
-Mean men $0.21$ $0.21$ $0.29$ $0.29$
-Quarter-District FE $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-Num. obs. $2,599$ $2,599$ $1,749$ $1,749$
-${}^{***}p<0.01$; ${}^{**}p<0.05$; ${}^{*}p<0.1$
-  • Notes: This table focuses on domestic violence cases and shows the association between a judge’s decision being appealed and the characteristics of the judges and of the decision.
-Columns (1) and (2) show the results for the sample of cases handled by final-stage judges.
-Columns (3) and (4) show the results for cases handled by early-stage judges.
-Table B.2 shows that female senior judges face appeal processes more frequently than male judges (Column (1)) when analyzing domestic violence cases.
-It also shows that this higher probability of facing an appeal process is partially explained by convicting more cases (Column (2)).
-Moreover, it shows that early-stage judges, whose conviction probabilities do not vary by gender according to Table 7, present appeal probabilities that do not vary by gender either (Columns (3) and (4)).
-Table B.3 presents qualitatively similar results for the probability of having a decision reversed.
-Table B.3: Association between Conviction Sentences and Reversal of Judges’ Decisions on Domestic Violence Cases
-Final Early
-Reversed Reversed Reversed Reversed
-(1) (2) (3) (4)
-Fem $0.02^{*}$ $0.02^{*}$ $-0.01$ $0.02$
-$(0.01)$ $(0.01)$ $(0.06)$ $(0.02)$
-Convicted $0.03^{***}$ $0.04^{**}$
-$(0.01)$ $(0.01)$
-Fem x Convicted $-0.01$ $-0.02$
-$(0.02)$ $(0.02)$
-Mean men $0.02$ $0.02$ $0.03$ $0.03$
-Quarter-District FE $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-Num. obs. $2,599$ $2,599$ $1,749$ $1,749$
-${}^{***}p<0.01$; ${}^{**}p<0.05$; ${}^{*}p<0.1$
-  • Notes: This table focuses on domestic violence cases and shows the association between the reversal of a judge’s decision and the characteristics of the judges and of the decision.
-Columns (1) and (2) show the results for the sample of cases handled by final-stage judges.
-Columns (3) and (4) show the results for cases handled by early-stage judges.
-We interpret these results as suggestive evidence that the gender conviction rate gap for domestic violence cases may have negative career consequences.
-For this reason, only female final-stage judges, who face weak career incentives, differ from their male counterparts when analyzing this type of case.
-Appendix C Modelling the Gender Conviction-Rate Gap
-In this appendix, we propose a simple economic model to formalize the concepts of the gender conviction-rate gap, in-group bias (Shayo and Zussman, 2011) and their driving forces, the representational and informational accounts (Boyd, Epstein, and Martin, 2010).
+In the end, our main outcome value equals $1$ if the case was classified into the category “defendant is guilty” and equals 00 otherwise.
+We also use the CPOPG dataset to measure recidivism (Section 6.2).
+We focus on whether the defendant recidivated within 2 years of their case’s final sentence.
+Following Possebom (2023), a defendant $i$ in a case $j$ recidivated ($Y_{ij}=1$) if and only if defendant $i$’s full name appears in a case $\bar{j}$ whose starting date is within 2 years after case $j$’s final sentence’s date.
+Importantly, case $\bar{j}$ can be about any type of crime, including very severe crimes, while case $j$ has to be about domestic violence, misdemeanor, or physical assault offenses.
+To match defendants’ names across cases, we use the Jaro–Winkler similarity metric and define a match if the similarity between full names in two different cases is greater than or equal to 0.95.181818Abramitzky2019 match full names in historical Censuses in the U.S.
+and Norway.
+They define a match between two individuals if the Jaro–Winkler similarity between their names is greater than or equal to 0.90 and if their dates of birth match exactly.
+Since we do not observe defendants’ dates of birth, we adopt a stricter Jaro-Winkler similarity threshold to define a match in our dataset.
+Moreover, in Table 8’s even columns, we exclude from the recidivism definition any case that is a domestic violence case.
+Specifically, we observe the offense type of case $\bar{j}$ and do not count it as a recidivism event if it is a domestic violence offense.
+Importantly, in Section 6.2 only, we delete the case-defendant pairs whose cases started in 2018 and 2019 because their recidivism variable is not properly defined due to right-censoring.
+Lastly, we use the CPOSG dataset to measure whether a case was analyzed by the Appeals Court and whether the Appeals Court’s ruling reverses the Trial Judge’s ruling.
+We start by merging the CJPG dataset with the CPOSG dataset using each case’s id code.
+When merging these datasets, we create an indicator variable that denotes which cases were analyzed by the Appeals Court, i.e., which cases were matched.
+This is the outcome variable in Table 6.1’s odd columns.
+We, then, randomly select 50 cases per year for manual classification (2010-2019) and divide them into three categories: “cases that went to the Appeals Court, but were immediately returned due to bureaucratic errors”, “cases whose trial judge’s rulings were affirmed” and “cases whose trial judge’s rulings were reversed”.
+Now, we use those 500 manually classified cases to train a classification algorithm. To do so, we divide them into a training sample (300 cases) and a validation sample (200 sentences).
+First, we design an algorithm to identify which cases went to the Appeals Court but were immediately returned. To do so, we simply check whether the Appeals Court’s decision is missing in our dataset.
+Finally, we design an algorithm to classify the non-empty cases into two categories: “cases whose trial judge’s sentences were affirmed” and “cases whose trial judge’s sentences were reversed”.
+To do so, we define a bag of words that were found to be strong signals of sentence reversal when we manually classified the cases in our sample.
+[^19] Possebom (2023, Appendix I.2) lists the entire bag of words.
+We, then, count how many times each one of those expressions appears in each sentence, and we normalize those counts to be between 0 and 1.
+Using the normalized counts, we train an L1-Regularized Logistic Regression using our training sample.
+We then validate this algorithm using our validation sample and find that it correctly classifies 96.2% of the cases.
+Given this high success rate, we use the L1-Regularized Logistic Regression algorithm to classify which trial judge’s rulings were reversed in our full sample.
+This is the outcome variable in Table 6.1’s even columns.
+Appendix B Modelling the Gender Conviction-Rate Gap
+In this appendix, we propose a simple economic model to formalize the concepts of gender conviction-rate gap and in-group bias (Shayo and Zussman, 2011), and their driving forces, the representational and informational accounts (Boyd, Epstein, and Martin, 2010).
 First, to formally define the first two concepts, we adapt the potential outcome model (Rubin, 1974) to a scenario where cases of different types of crime may be treated with a female judge instead of a male judge.
-Second, to formally define the representational and informational accounts of the in-group bias, we adapt the threshold-crossing model (Heckman and Vytlacil, 2005) to a context where the latent heterogeneity is associated with the evidence of a criminal case and the propensity score is associated with the judge’s conviction criteria.
+Second, to formally define the representational and informational accounts of the in-group bias, we adapt the threshold-crossing model (Heckman and Vytlacil, 2005) to a context where the latent heterogeneity is associated with the evidence of a criminal case and the propensity score function is associated with the judge’s conviction criteria.
 We start by looking at criminal cases as our observational unit.
@@
 We define the gender conviction-rate gap as
-$\mathbb{E}\left[Y\left(1,1\right)-Y\left(0,1\right)\right]\neq 0,$ (C.1)
+$\mathbb{E}\left[Y\left(1,1\right)-Y\left(0,1\right)\right]\neq 0,$ (B.1)
 i.e., the probability that a domestic violence case ends in a conviction when presided by a female judge is different from the probability that a domestic violence case ends in a conviction when presided by a male judge.
 We also define the in-group bias as
-$\left\{\begin{matrix}[l]\mathbb{E}\left[Y\left(1,1\right)-Y\left(0,1\right)\right]&>&0\\ \mathbb{E}\left[Y\left(1,1\right)-Y\left(0,1\right)\right]&>&\mathbb{E}\left[Y\left(1,0\right)-Y\left(0,0\right)\right]\end{matrix}\right.,$ (C.2)
+$\left\{\begin{matrix}[l]\mathbb{E}\left[Y\left(1,1\right)-Y\left(0,1\right)\right]&>&0\\ \mathbb{E}\left[Y\left(1,1\right)-Y\left(0,1\right)\right]&>&\mathbb{E}\left[Y\left(1,0\right)-Y\left(0,0\right)\right]\end{matrix}\right.,$ (B.2)
 i.e., the gender conviction-rate gap for domestic violence cases (i) is positive, and (ii) is larger than the gender conviction-rate gap for other types of crime.
@@
 The judge convicts the defendant if the judge interprets the evidence as being sufficient to do so, i.e.,
-$Y=\mathbf{1}\left\{P\left(F,D,W\right)\leq E\left(F,D,R,V\right)\right\},$ (C.3)
+$Y=\mathbf{1}\left\{P\left(F,D,W\right)\leq E\left(F,D,R,V\right)\right\},$ (B.3)
 where $Y$ denotes whether the defendant is convicted or not.
-Note that Equation (C.3) imposes two exclusion restrictions.
+Note that Equation (B.3) imposes two exclusion restrictions.
 While the relationship status between the victim and the defendant — $W$ —- only enters the judge’s conviction criterion function, being caught red-handed — $R$ — only enters the evidence interpretation function.
@@
 Using this model, we can formally define the representational account driving the in-group bias as
-$P\left(1,1,1\right)-P\left(0,1,1\right)<P\left(1,1,0\right)-P\left(0,1,0\right)\leq 0,$ (C.4)
+$P\left(1,1,1\right)-P\left(0,1,1\right)<P\left(1,1,0\right)-P\left(0,1,0\right)\leq 0,$ (B.4)
 i.e., female judges may set stricter conviction criteria than male judges for domestic violence cases and female judges are even stricter when the case is associated with intimate partner violence.
 We can also formally define the informational account driving the in-group bias as
-$E\left(1,1,0,v\right)-E\left(0,1,0,v\right)>E\left(1,1,1,v\right)-E\left(0,1,1,v\right)\geq 0,$ (C.5)
+$E\left(1,1,0,v\right)-E\left(0,1,0,v\right)>E\left(1,1,1,v\right)-E\left(0,1,1,v\right)\geq 0,$ (B.5)
 for any $v$ in the support of $V$.
@@
 In the main text, we use data from criminal offenses in the State of Sao Paulo, Brazil, to test the existence of each one of the four concepts defined above.
-In Table 3, we show that the gender conviction-rate gap (Equation (C.1)) exists.
-Then, in Table 4, we show that in-group bias (Equation (C.2)) is present in our application.
-Finally, Table 5 shows that our data is consistent with the representational account (Equation (C.4) while Table 6 provides evidence that our data is consistent with the informational account (Equation (C.5).
+In Table 3, we show that the gender conviction-rate gap (Equation (B.1)) exists.
+Then, in Table 4, we show that in-group bias (Equation (B.2)) is present in our application.
+Finally, Table 5 shows that our data is consistent with the representational account (Equation (B.4) while Table 6 provides evidence that our data is consistent with the informational account (Equation (B.5).